### v2

# DIAS: A Comprehensive Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

Abstract
Digital subtraction angiography (DSA) is universally acknowledged as the gold standard for examining lesion angioarchitecture, elucidating arterial blood supply dynamics, and guiding endovascular interventions. The automatic segmentation of intracranial arteries (IA) in DSA, which is pivotal for quantifying vascular morphology, plays an essential role in computer-assisted stroke research and clinical practices. Nevertheless, research in this specific domain remains constrained, primarily owing to the unavailability of publicly datasets for IA segmentation within the research community. Currently, the predominant focus of methodologies lies in the segmentation of single-frame DSA using in-house datasets. These methods, limited by the partial inclusion of contrast in single-frame DSA, encounters challenges in rendering a precise representation of vascular structures. In this paper, we introduces DIAS, a dataset specifically developed for IA segmentation in DSA sequences. A comprehensive benchmark has been established for evaluating DIAS, covering fully, weakly, and semi-supervised segmentation methods. Specifically, we propose a vessel sequence segmentation network that captures the spatiotemporal representation of intravascular contrast for segmenting vessels in DSA sequences. For weakly-supervised learning, we propose a novel scribble learning-based image segmentation framework, incorporating both scribble supervision and consistency regularization. Furthermore, we introduce a random patch-based self-training framework that harnesses unlabeled DSA sequences to improve segmentation performance. Our extensive experiments on the DIAS dataset demonstrate the effectiveness of these methods as potential baselines for future research and clinical applications. The dataset and code will be publicly available at https://github.com/lseventeen/DIAS.
Index Terms:
digital subtraction angiography, intracranial artery segmentation, dataset, semi-supervision learning, weakly supervision learning
I Introduction
Cerebrovascular diseases significantly contribute to global mortality and long-term disability [1]. In recent decades, there has been a notable decline in the mortality rate associated with intracranial artery (IA) diseases, such as intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, largely attributed to rapid advancements in medical and imaging technology [2]. Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3]. It removes bony structures by subtracting the pre-contrast image, thus preserving the vessels filled with contrast agent. Owing to its inherently superior spatial and temporal resolution, DSA can accurately elucidate lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2]. Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
Conventionally, neuroradiologists undertake the visual analysis of DSA for diagnostic purposes, clinical decision-making, and postoperative assessment, which is inherently time-consuming, labor-intensive, and subjective. Automated segmentation of vessels DSA harbors the potential to mitigate this challenge by delineating vascular morphology. Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7]. Despite the undeniable significance of cerebrovascular segmentation in DSA, it remains an under-researched area. One possible reason is the current absence of a DSA dataset tailored specifically for the formulation and evaluation of cerebrovascular segmentation techniques in the medical imaging community. However, given the intrinsic attributes of its imaging modality, the task of creating a brain DSA vascular segmentation dataset is immensely challenging. Primarily, the DSA sequence presents a projection of intravascular contrast media from a consistent viewing angle, thereby visualizing only a fraction of the IA within the contrast agent in any given image, as illustrated in Fig. 1. In the process of annotating vessels within DSA, neuroradiologists are required to conduct a detailed examination of each image frame, which is both laborious and expensive [8]. Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures in imagery, diminished contrast, and multifaceted background disturbances, accentuates the difficulties associated with annotation.
Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8]. Yet, all these efforts are anchored to proprietary datasets, absent a consistent evaluation benchmark. Furthermore, most of these menthods [9, 10] are limited to single-frame image segmentation, overlooking the rich temporal characteristics inherent in DSA sequences. Segmentation methods based on single-frame DSA may not comprehensively capture the vascular intricacies of IA, potentially limiting their efficacy in diagnosing and treating cerebrovascular diseases [8]. In DSA sequences, vascular segmentation is characterized as a dimensionality-reduction task, taking a 2D+time image input to yield a 2D mask output. This task deviates from traditional 2D and 3D vascular segmentation meathods [11, 12, 13]. The pivotal challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
Figure 1: A representative sample of intracranial vessels in sagittal view from DSA, where t to t+5 are consecutive frames of the arterial phase.
In this work, we construct the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons. Given the intricate and labor-intensive nature of vessel annotation within DSA sequences, we’ve provided two types of weak annotations for training samples, aiming to facilitate studies on weakly-supervised IA segmentation. For the DIAS, we have devised three benchmarks tailored for fully-supervised segmentation (FSS), weakly-supervised segmentation (WSS), and semi-supervised segmentation (SSS): the vessel sequence segmentation network, the scribble learning-based segmentation framework and the random patch-based self-training framework. We have carried out comprehensive experiments using the DIAS, establishing robust baselines that will serve as a foundation for future algorithmic comparisons on the DIAS dataset. We hope that the DIAS dataset could facilitate open research in cerebral vessel segmentation.
TABLE I: Summary of several public vessel segmentation datasets. ${}^{\dagger}$ represents only training set data is available.
Dataset Year Target Modality Dimension Numbers Resolution
DRIVE [14] 2004 retinal vessels color fundus 2D 40 565×584
CHASE_DB1 [15] 2012 retinal vessels color fundus 2D 28 999$\times$960
STARE [16] 2000 retinal vessels color fundus 2D 20 700$\times$605
HRF [17] 2013 retinal vessels color fundus 2D 45 3504$\times$2336
ROSE1 [18] 2020 retinal vessels OCTA 2D 117 304$\times$304
ROSE2 [18] 2020 retinal vessels OCTA 2D 112 512$\times$512
CHUAC [19] 2018 coronary artery angiograpy 2D 30 189$\times$189
DCA1 [20] 2019 coronary artery angiograpy 2D 134 300$\times$300
${}^{\dagger}$ [21] 2022 coronary artery CTA 3D 40 512$\times$512$\times$200
${}^{\dagger}$ [22] 2014 lung vessel CT 3D 20 -
DIAS 2023 intracranial artery DSA 2D+Time 120 (753) 800$\times$800$\times$4-14
II Related work
II-A Vessel Segmentation Datasets
Vessel segmentation is a popular area of research within computer-aided medicine, which has witnessed remarkable progress in recent years, owing to technological advancements and the availability of high-quality annotated datasets. We have summarized the several publicly available vessel segmentation datasets in Table I. A representative example is retinal vessel segmentation, which has many publicly available datasets due to its easy availability, such as colour fundus datasets (DRIVE [14], CHASE_DB1 [15], STARE [16], HRF [17]) and OCTA datasets (ROSE1 and ROSE2 [18]). These datasets consist of 20-117 2D images with spatial resolution ranging from 304$\times$304 to 3504$\times$2336. However, other vessel segmentation datasets are relatively scarce [8]. CHUAC [19] and DCA1 [20] are coronary angiography datasets with resolutions of 189$\times$189 and 300$\times$300, respectively. ASOCA [21] and VESSEL12 [22] are the 3D vessel segmentation datasets of coronary artery CTA and lung CT from the challenges, in which only training set data is available.
The availability of such datasets has led to a surge of outstanding works in the field of segmentation, with researchers striving to develop algorithms that outperform existing methods. So far, there have been no publicly available DSA datasets with vessel annotation in the medical imaging community, and only a few studies conducted with in-house datasets. DCVessel [9] comprises 30 DSA arterial-phase images at a resolution of 700$\times$700, while HEART [23] comprises 1092 DSA coronary images at 256$\times$256 resolution. Su et al. [8] manually segmented arteries and veins on 97 DSA series, covering both anterior-posterior and lateral views, each with a resolution of 1024$\times$1024 pixels. In this work, we have released DIAS, an IA segmentation dataset derived from DSA sequences. It comprises 120 sequences (753 frames), each with a resolution of 800$\times$800 pixels, and the lengths of these sequences range from 4 to 14. The DIAS can serve as a reliable benchmark for the development of algorithms on IA segmentation from DSA sequences. We hope that this dataset will intrigue researchers to explore this field further.
II-B Vessel Segmentation Methods
In the past few years, deep learning has demonstrated great potential in vessel segmentation, surpassing traditional methods, such as hand-crafted features [24], filtering-based models [25], and statistical models [26]. Deep learning-based methods have become primary approaches for vessel segmentation, with UNet [27] serving a crucial role as a milestone network architecture, giving rise to numerous variant networks. These networks utilize an encoder-decoder architecture with popular components, such as multi-scale design [13, 9, 28], deep supervision [29, 9, 12], attention mechanisms [30, 13, 31], atrous convolution [32, 29, 12, 9], and improved skip connections [9, 12]. Moreover, several novel neural networks [33, 34] have also succeeded in vessel segmentation by leveraging generative adversarial networks (GAN) or graph neural networks (GNN). Some studies focused on vessel segmentation beyond network architecture, such as optimized training strategies [35, 36] and modifying loss functions [37].
However, there have been few studies on DSA vessel segmentation. Meng et al. [9] proposed a CNN-based multiscale dense CNN segmentation framework, which includes a multiscale atrous convolution module, an improved dense block, and redesigned skip connections, for IA segmentation in single frame DSA. Zhang et al. [23] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation. These two approaches are tailored for vessel segmentation in individual DSA frames. For vessel sequence, Hao et al. proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features [38]. Su et al. proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8]. While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
Fully annotating vessel datasets is both time-consuming and costly. Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers. WSL trains models with sparse annotations instead of full ones. The common methods involves online labeling [39] via consistency checks with the mean teacher or using pseudo-labels supervision [40] derived from model predictions. SSS aims to explore tremendous unlabeled data with supervision from limited labeled data. Recently, self-training methods [41, 42] have dominated this field. Furthermore, methods employing consistency regularization strategies [41, 43, 44] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image. To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences. In this regard, we propose a comprehensive benchmark involving FSS, WSS, and SSS for IA segmentation based on the constructed DSA-sequences dataset.
Figure 2: Full annotated process of DIAS.
Figure 3: Full annotated representative samples of DIAS at coronal view and sagittal view.
Figure 4: Examples of two forms scribble annotations: the scribbles annotation with little clinical experience (SALE) and Random Drawing based the Full Annotation (RDFA). Red represents annotated vessel pixels, blue represents annotated background pixels.
III DIAS Dataset
III-A Dataset Description
The DIAS dataset was derived from intraoperative DSA images obtained during ischemic stroke treatments, including ICAS and MCAO, at Beijing Tiantan Hospital. The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study. All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details. The original DSA sequences were captured by the angiography machine during the actual endovascular therapy process and stored as DICOM files. It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view, lateral view or both. In this study, we retained 120 sequences after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates. The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS). The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt. We retained all the arterial-phase frames and removed the pre-contrast, capillary-phase, and venous-phase frames under the guidance of a neurosurgeon. As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
III-B Reference Annotation Protocol
III-B1 Full Annotation
Figure 5: Illustration of the proposed pipeline for DSA-sequence intracranial artery segmentation. (a) Vessel sequence segmentation network with (b) sequence feature extraction module; (c) WSL: scribble learning-based segmentation composed of scribble supervision and consistency regularization; (d) SSL: random patch-based self-training framework. (WSC: Weight Sharing Convolution, SF: Sequence Fusion, GMP: Global Max Pooling, WSD: Weight Sharing Downsampling, DA: Data augmentation, WA: Weak Augmentation, SA: Strong Augmentation)
Manual annotation of IA in DSA sequences requires taking into account all the image information in the sequence. Fig. 1 shows an example of intracranial vessels in five consecutive frames of the arterial phase in the sagittal view. During this period, the contrast agent is gradually dispersed from the internal carotid artery (ICA) to cerebral vessels such as the anterior cerebral artery (ACA), middle cerebral artery (MCA), and posterior cerebral artery (PCA) with blood flow. For a single image, the contrast agent is visible only in a portion of the IA. For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+4, it filled the ACA, MCA, and PCA. Annotating the IA of each image in the sequence and then merging the annotation mask can maximize the utilization of the sequence’s vessel information to obtain accurate annotations. The thinness of vessels, particularly in later-stage arterial images, renders the commonly used strategy of outlining their borders in medical image annotation unfeasible. Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure. Fig. 3 shows that our proposal for a compromise solution, which reduces workload while maintaining annotation accuracy:
  • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the later phase. The contrast agent during the anterior phase predominantly fills the ICA as opposed to the smaller vessels such as the ACA, MCA, and PCA arteries, which are filled by the contrast agent in the later phase.
  • The last image of each phase is annotated pixel-wise, followed by a meticulous review of other same-phase images to finalize the annotation. Thick vessels are the focus of annotation in the anterior phase, while thin vessels are prioritized in the later phase.
  • The annotations of thick and thin vessels are merged using a pixel-level OR operation (1 for vessel pixel and 0 for background) to create the ultimate vessel mask.
We randomly selected 60 sequences from DIAS for the purpose of annotating blood vessels, using the patient ID as the criterion for selection. The remaining 60 sequences were used for research on IA semi-supervised segmentation. The DSA sequence was initially annotated by two MD students of surgery, followed by corrections from an associate physician with 10 years of experience, and finally confirmed and double-checked by a chief physician with 15 years of experience.
III-B2 Scribble Annotation
Scribble-based image segmentation has proven to be an efficacious and sought-after method for WSS [45]. This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries. To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based the Full Annotation (RDFA). SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig. 4 (a). Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging. Consequently, pixels in these regions were annotated during annotation. Moreover, scribble segmentation relies on supervised information for every category. Apart from annotating the vessels, we conducted extensive scribble annotations for the background wherever possible, as illustrated by the blue marker in Fig. 4 (a). Follow [46], we generate RDFA by skeletonizing and randomly pruning each category of the full annotation, as shown in Fig. 4 (b). RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations. In SALE and RDFA, we set the pixel values of the background and vessel annotations to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.
IV Methods
Automatic IA segmentation in DSA sequences is an under-explored and challenging task. To fully explore potential research topics, three elaborately designed benchmarks for IA segmentation were proposed, aiming to provide comprehensive benchmarks for FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
IV-A Fully-supervised Segmentation
Segmenting IA in DSA sequence can be considered as a dimensionality reduction (DR) task. Given an image $I\in\mathbb{R}^{\prod_{d=1}^{N}n_{d}}$, where $N$ represents the number of dimensions and $n_{d}$ is the size of the image in the corresponding dimension $d$, DR segmentation can be defined as finding a function $f$ such that $f:I\rightarrow O$, where $O\in\mathbb{R}^{\prod_{d=1}^{M}n_{d}}$ represents segmentation output and $M$ ($<N$) is its dimensionality [47]. For IA segmentation in DSA sequence, $N=3$, $M=2$. Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representation of vessels from various scales to achieve precise IA segmentation in DSA. Figure 5 (a,b) illustrates an overview of the proposed network architecture. VSS-Net is composed of multi-scale sequence feature extraction modules (SFEM). These modules handle the 2D+T input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations. Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
In the encoder, the input feature map from each stage is integrated with the raw sequence, serving to supplement fine-grained semantic information, except for the initial stage. Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation. Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1x1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels. This process is formulated as:
$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i})$ (1)
where, $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence. $W$ are $1\times 1$ convolutional kernels. $\ast$ denotes the convolution operation. $\oplus$ denotes the concatenation operation.
Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of sequence. ConvGRU [48] is an extension of the traditional GRU that integrates convolutional operations, making it more suitable for spatial data like images. Consider $x_{t}$ as the input to the ConvGRU, with corresponding hidden states denoted by $h_{t}$, we can delineate the updating process of ConvGRU unrolled by time as follows:
$z_{t}$ $=\sigma(W_{z}\ast x_{z}+U_{z}\ast h_{t-1})$ (2)
$r_{t}$ $=\sigma(W_{r}\ast x_{r}+U_{r}\ast h_{t-1})$ (3)
$\tilde{h}_{t}$ $=\tanh(W_{h}\ast x_{t}+U_{h}\ast(r_{t}\odot h_{t-1}))$ (4)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (5)
where, $\sigma$ represents the sigmoid function. $\tanh$ represents the hyperbolic tangent function. $\odot$ denotes the Hadamard product, i.e., element-wise multiplication. $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
We simultaneously capture the spatiotemporal dynamics of vascular flow in both forward and backward directions. Subsequently, the outputs of the bidirectional ConvGRU are fused into a single sequence [49]. The resulting integrated features encompass angiographic representations of both macrovascular and microvascular structures, spanning the early to late arterial stages. The process can be formulated as follows:
$h_{t}^{f}$ $=ConvGRU(h_{t-1},x_{t})$ (6)
$h_{t}^{b}$ $=ConvGRU(h_{t+1},x_{t})$ (7)
$h_{t}$ $=\tanh(W_{f}\ast h_{t}^{f}+W_{b}\ast h_{t}^{b})$ (8)
where, $h_{t}^{f}$ and $h_{t}^{b}$ denote the hidden states from the forward and backward ConvGRU units, respectively. $x_{t}$ is the $t^{th}$ output feature from WSC.
The output feature maps are spatially downsampled using stride convolution with weight sharing to serve as inputs for the next stage encoder. Simultaneously, we employ global max pooling along the sequence dimension of the feature maps for dimensionality reduction. The generated multi-scale feature maps are compatible with any U-shaped segmentation architecture. In particular, we integrate it with FR-UNet, where the outputs of the multi-scale SFEM are fed into each stage of FR-UNet. The resultant VSS-Net effectively harnesses the strengths of both components, extracting spatiotemporal representations to realize precise vessel segmentation within sequences.
IV-B Weakly-supervised Segmentation
A common strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled points. However, this baseline approach does not utilize the unlabeled points that contain critical semantic information. The performance degradation is unavoidable due to the lack of labeled data points. To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels. To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization (CR). Fig. 5 (b) presents the implementation framework for SSCR. SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
We define $P$ as the set of image pixels and $S\subseteq P$ as the set of annotated pixels. The proposed approach feeds $P$ into two segmentation networks with different augmentation strategies. The outputs of each network are supervised by applying the cross-entropy function solely to the annotated pixels on the scribble labels. The function, namely partial cross entropy (pCE), can be formulated as:
$\mathcal{L}_{pce}(y,\hat{y})=-\sum_{i=1}^{M}\sum_{j=1}^{N}{\mathbbm{1}(p_{i,j}\in S)(y_{i,j}\log(\hat{y}_{i,j|\theta})})$ (9)
where, $M$ and $N$ represent the width and height of the image, respectively. $\hat{y}_{i,j|\theta}$ is the predict for the pixel $p_{i,j}\in S$ given the network parameters $\theta$, and $y_{i,j}$ denotes the ground truth label.
To generate pseudo labels, we first calculate the class distribution of the one-hot prediction map using limited softmax, which exclusively calculates the channel corresponding to the annotated class. Then, we apply scribble annotations to correct the pseudo labels in order to improve their confidence. The formula for creating the pseudo label is as follows:
$PL_{i}=\begin{cases}s_{i,j},&\text{if}\ p_{i,j}\in S\\ \text{argmax}(\text{softmax}(\mathcal{H}(p_{i,j};\theta)[C])),&\text{otherwise}\par\end{cases}$ (10)
where, $s_{i,j}\in S$ defines annotated pixels, and $\mathcal{H}(p_{i,j};\theta)$ defines the predicted output of $p_{i,j}$. $C$ is the channel of annotated class, i.e., 0 and 1.
We enforce network consistency by utilizing cross pseudo supervision, which employs pseudo labels from one network to supervise the other network [43]. Additionally, these pseudo labels are imposed a consistency loss to penalize inconsistent segmentation. The training objective contains three losses: scribble supervision loss $\mathcal{L}_{scr}$, cross pseudo supervision loss $\mathcal{L}_{cps}$, and consistency loss $\mathcal{L}_{con}$:
$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL))$ (11)
$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A}))$ (12)
$\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B})$ (13)
where, $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function. $P_{A}$ and $P_{B}$ are the outputs of two models. $PL_{A}$ and $PL_{B}$ are the pseudo labels. $SL$ is the scribble label.
Finally, the training objective $\mathcal{L}$ is formulated as
$\mathcal{L}=\mathcal{L}_{scr}+\lambda_{1}\mathcal{L}_{cps}+\lambda_{2}\mathcal{L}_{con}$ (14)
where, $\lambda_{1}$ and $\lambda_{2}$ are the weight of loss components $\mathcal{L}_{cps}$ and $\mathcal{L}_{con}$.
IV-C Semi-supervised Segmentation
Annotation of cerebral vessels manually from DSA image sequences is a laborious and time-consuming task. Semi-supervised learning has shown the potential to alleviate this problem by learning from a large set of unlabeled images and limited labeled samples [42]. The primary challenge in semi-supervised setting lies in how to effectively utilize the unlabeled images. Self-training [50] via pseudo labeling is a conventional, simple, and popular pipeline to leverage unlabeled data, where the retrained student is supervised with hard labels produced by the teacher trained on labeled data, which is commonly regarded as a form of entropy minimization in semi-supervised learning. For semi-supervised segmentation problem, let $\mathcal{D}^{l}=\{(x_{i},y_{i})\}_{i=1}^{M}$ be pixel-wise labeled images and $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$ be unlabeled images, where in most cases $N\gg M$. Self-training can be simplified into the following steps:
  1. [leftmargin=*,itemsep=-5pt]
  2. 1.
    Train a teacher model $T$ on labeled data $\mathcal{D}^{l}$;
  3. 2.
    Predict hard pseudo labels on unlabeled data $\mathcal{D}^{u}$ with $T$ to obtain $\mathcal{\hat{D}}^{u}=\{(u_{i},T(u_{i})\}_{i=1}^{N}$;
  4. 3.
    Re-train a student model $S$ on the label data and pseudo-label data $\mathcal{D}^{l}\cup\mathcal{\hat{D}}^{u}$ for final evaluation.
We follow the plainest self-training strategy to provide a SSS benchmark for DIAS. The injection of strong data augmentations on unlabeled images is highly advantageous as it decouples the predictions as well as alleviates overfitting on noisy pseudo-labels. Despite its simplicity, self-training with strong data augmentations surpasses existing methods without any bells and whistles [41]. We performed two types of data augmentations in our implementation: weak and strong, denoted by $A^{s}$ and $A^{w}$ respectively. Weak data augmentations (WDA) consist of horizontal flipping, vertical flipping, and $[90,180,270]$ degree rotation. Strong data augmentations (SDA) include Cutout, adding white Gaussian noise, Gaussian blurring, and elastic deformation. We randomly adjust the contrast of each image in the DSA sequence to simulate angiography with intensity differences. We randomly extract patches from the DSA sequence to train the teacher and student models. Notably, the patches in each mini-batch were randomly extracted from $\mathcal{D}^{l}$ and $\mathcal{\hat{D}}^{u}$ with a probability $P$ when retraining the student model. Additionally, the final performance can be further improved by switching the teacher-student role and relabeling unlabeled images for iterative training. The pseudocode of the random patch-based self-training framework, referred to as RPST, is provided in Algorithm 1.
Input: Labeled training set $\mathcal{D}^{l}=\{(x_{i},y_{i})\}_{i=1}^{M}$,
          Unlabeled training set $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$,
          Weak/strong augmentation $\mathcal{A}^{w}/{A}^{s}$,
          Teacher/student model $T/S$,
          Execution probability $P\in(0,1)$
Output: Trained student model $S$
Train $T$ on $\mathcal{D}^{l}$ with loss $\mathcal{L}$
Obtain pseudo labeled $\mathcal{\hat{D}}^{u}=\{(u_{i},T(u_{i}))\}_{i=1}^{N}$
for Iterative training $N$  do
       for Minibatch with the size of $B$  do
             for $k\in\{1,...,B\}$ do
                   if  Random number $r>P$ then
                         Randomly extract patch $(x^{p}_{k},y^{p}_{k})$ from $\mathcal{D}^{l}$
                         $x^{p}_{k},y^{p}_{k}\leftarrow(\mathcal{A}^{s}\&\mathcal{A}^{w})(x^{p}_{k},y^{p}_{k})$
                   else
                         Randomly extract patch $(x^{p}_{k},y^{p}_{k})$ from $\mathcal{\hat{D}}^{u}$
                         $x^{p}_{k},y^{p}_{k}\leftarrow\mathcal{A}^{w}(x^{p}_{k},y^{p}_{k})$
             end for
            $\hat{y}^{p}_{k}=S(x^{p}_{k})$
             Update $S$ to minimize $\mathcal{L}$ of $\{(\hat{y}^{p}_{i},y^{p}_{k})\}_{i=1}^{B}$
       end for
      if Not the last iterative then
             Obtain new pseudo labeled $\mathcal{\hat{D}}^{u}=\{(u_{i},S(u_{i}))\}_{i=1}^{N}$
end for
Algorithm 1 RPST
V Experiments
V-A Implementation Details and Evaluation Metrics
The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level. To facilitate input into the model, all DSA sequences were resampled to a length of 8. All DSA sequences were z-score normalized. Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64. The model parameters were optimized using AdamW [51] with an initial learning rate of 5e-4. The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs. The parameter $P$ was set to 0.5, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment. For SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5 . We only use one network to generate the results for evaluation. All experiments used PyTorch with constant hyperparameters and were conducted on a single GeForce RTX 3090 GPU. We calculated the dice similarity coefficient (DSC), the area under the receiver operating characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and intersection over union (IOU) to evaluate the performance of DSA sequence segmentation. Moreover, we introduced the vascular connectivity (VC) [12] metric, which calculates the ratio of connected components in the segmentation image to those in the ground truth, as a means of measuring the connectivity of vessels.
V-B Fully-supervised Segmentation
[…]
$y=\mathcal{H}_{2d}(Squ(x))$ (15)
where, $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
For a 3D model $\mathcal{H}_{3d}$, the DSA sequence is fed directly into the model, where it is treated as volume data. The dimension, $C$, of the model output is squeezed by DR operation. Finally, $1\times 1$ convolution reduces $F$ to the number of classes:
$y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x)))$ (16)
Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods. Using FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe. Specifically, Sen has increased by 1.3%. ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe. However, the number of model parameters is approximately 5.58 times that of VSS-Net. The VC score that is 19.57 higher than VSS-Net indicates poorer vascular connectivity. Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages. PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures. Upon overall comparison of the DSC scores for all ‘DR+2D model’ and ‘DR+3D model’, 2D models slightly outperform 3D models with fewer parameters. Although in terms of data format, 2D+time DSA sequences are equivalent to 3D data, they do not have a one-to-one spatial relationship with the ground truth. More importantly, 3D convolutions can only model local temporal representations. The ‘DR+2D model’ benefits from treating ‘sequences’ as ‘channels’, which enables it to learn global representations along the sequence dimension while performing 2D convolutions, which aggregates contrast agent responses that reflect vascular morphological information in the sequence direction. However, the approach of setting fixed channel numbers is not applicable to variable image sequences. We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Figure 6. In the figure, the red and blue colors represent pixels that are false negatives and false positives, respectively, indicative of segmentation errors. It can be observed that a prevalent issue shared by all methods is the abundant incorrect segmentation of thin vessels. Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Figure 6. In comparison with other methods, VSS-Net has the highest number of white pixels in the fine vessels, indicating the best segmentation results.
TABLE II: Results of fully-supervised segmentation with state-of-the-art models.
Methods Params / M Metrics
[…]
2D UNet [27] 31.04 0.7677 0.9647 0.757 0.9829 0.6258 0.9811 68.29
Res-UNet [54] 9.37 0.7656 0.9655 0.7425 0.985 0.6235 0.9816 60.56
UNet++ [53] 9.05 0.7768 0.9659 0.7726 0.983 0.6374 0.983 54.65
Att-UNet [52] 8.73 0.7760 0.9660 0.7686 0.9835 0.6258 0.9811 59.37
[…]
MAA-Net [13] 5.80 0.7693 0.9661 0.7423 0.9855 0.6361 0.9818 68.18
FR-UNet [12] 7.37 0.7779 0.9665 0.7646 0.9841 0.6376 0.9827 53.92
3D UNet [27] 90.27 0.7691 0.9648 0.769 0.9824 0.6275 0.9819 62.92
Res-UNet [54] 26.38 0.7701 0.9653 0.7597 0.9835 0.6284 0.9817 73.05
UNet++ [53] 26.17 0.7649 0.9639 0.7711 0.9811 0.6214 0.9811 67.02
Att-UNet [52] 23.6 0.7677 0.9647 0.757 0.9829 0.6258 0.9811 68.29
[…]
FR-UNet [12] 20.57 0.7635 0.9646 0.7518 0.983 0.621 0.9815 65.67
[…]
PSC [47] 6.43 0.7649 0.9646 0.7552 0.9832 0.6223 0.9814 76.04
SVS-Net [38] 8.48 0.7698 0.9659 0.7474 0.9853 0.6277 0.9803 97.28
ST-UNet [8] 104.37 0.7818 0.9675 0.7533 0.9864 0.6436 0.9843 62.59
VSS-Net 18.69 0.7822 0.9668 0.777 0.9838 0.6442 0.9857 43.02
Figure 6: Visualization of segmentation results of state-of-the-art models. The first row is coronal view, and the second row is sagittal view. The red pixel indicates false negatives, and the green pixel indicates false positives.
TABLE III: Results of Scribble-supervised IA sequence segmentation.
Methods RDFA SALE
[…]
pCE [47] 0.7455 0.9586 0.7845 0.9735 0.5966 0.9379 60.11 0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
EM [56] 0.7458 0.9575 0.8029 0.9707 0.5970 0.9725 41.18 0.6356 0.9226 0.8508 0.9281 0.4699 0.9618 9.05
RLoss [45] 0.7146 0.9600 0.6539 0.9867 0.5581 0.9367 46.30 0.5916 0.9147 0.7812 0.9257 0.4232 0.9475 5.17
IVM [46] 0.6606 0.9299 0.8546 0.9360 0.4971 0.9455 71.58 0.6322 0.9198 0.8760 0.9232 0.4662 0.9628 14.78
DMPLS [40] 0.7471 0.9581 0.7963 0.9717 0.5995 0.9708 37.56 0.6327 0.9208 0.8508 0.9259 0.4668 0.9643 9.86
SSCR 0.7589 0.9603 0.8079 0.9732 0.6141 0.9761 36.52 0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
Figure 7: Visualization of the segmentation results of scribble supervised segmentation methods trained on RDFA and SALE annotations.
V-C Weakly-supervised Segmentation
We compared our method with five scribble-supervised segmentation methods: partial Cross Entropy (pCE) [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40]. To ensure a fair comparison, all experimental networks employed the UNet [27], consistent with the approach in DMPLS. Table III presents the results of scribble-supervised segmentation of the DIAS dataset. Fig. 7 showcases the visual results of all methods. Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods. It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 4.67% higher compared to SSCR, and it is 4.12% higher on SALE. Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background. The visualization results from IVM indicate that achieving high Sen comes at the cost of introducing numerous incorrectly segmented vessel pixels. Consequently, this leads to lower scores in other evaluation metrics such as the DSC and AUC. Regarding SALE, the vessels segmented by the same method are thicker than the ones trained with RDFA annotation, leading to a larger number of false positive pixels at the vessel’s edge. This outcome may have been caused by the SALE focusing on the annotation of thick vessels. Similarly, due to the lack of supervision for small vessels, where vascular connectivity issues are more prevalent, SALE struggles to extract fine vessels, resulting in a VC score significantly better than RDFA. From the last column of the figure, it is evident that SSCR’s segmentation results have the least number of blue pixels, corroborating the best quantitative results presented in the Table III.
TABLE IV: Ablation studies results of SSCR for IA sequence segmentatio with SALE annotation. (pCE: partial Cross Entropy, PS: Pseudo-labeling Supervision, CPS: Cross Pseudo Supervision, CR: Consistency Regularization)
Methods Metrics
[…]
0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
0.6364 0.9244 0.8346 0.9316 0.4706 0.9479 6.57
0.6472 0.9275 0.8385 0.9346 0.4827 0.9445 7.29
0.6484 0.9276 0.8536 0.9334 0.4827 0.9639 8.28
0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
TABLE V: Ablation research results of RPST for semi-supervised IA sequence segmentation.
Methods Data Metrics
[…]
Teacher 1 0 0.6266 0.9259 0.6853 0.9451 0.4693 0.8538 385.62
Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 107.35
Student w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 108.96
Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 103.09
Teacher 3 0 0.6919 0.949 0.7358 0.9673 0.5336 0.9182 320.02
Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 121.84
Student w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.57 0.9578 125.83
Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 125.78
Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.591 0.9694 148.89
Student 10 30 0.7544 0.9618 0.766 0.9786 0.6087 0.9769 93.31
Student w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
Student 10 60 0.7579 0.9625 0.7662 0.9795 0.613 0.9786 87.72
We conducted ablation studies to investigate the impact of various components in our proposed pipeline for weakly-supervised IA segmentation. As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis. Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as DSC and AUC. Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC. Compared to a single-model approach, a dual-model setup, receiving the same scribble supervision, generates pseudo-labels for mutual supervision. This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability. We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation without CPS. Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC. Additionally, it has attained the highest Sen. We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models. The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 2.5% and 2.38%, respectively.
V-D Semi-supervised Segmentation
The number of labeled and unlabeled samples and strong data augmentation (SDA) affect the performance of RPST. In this section, we randomly select 1, 3, and 10 labeled DSA sequences from the DIAS training set for SSS experiments. We leverage 30 and 60 DSA unlabeled sequences to generate pseudo labels and participate in training student. Table V shows that pseudo-labeled data generated by the trained teacher significantly improves the performance of the model, especially when there are fewer labeled data available. For example, the DSC of the student is improved by 2.2% with one labeled data and 30 unlabeled data compared to the teacher. This performance gain is positively correlated with the amount of unlabeled data. Ablation studies on SDA demonstrate the effectiveness of enforcing consistency through the application of both strong and weak data augmentation. While the pseudo-labeled data used to train the student model contains noisy labels, they can still enhance the model’s generalization capability, albeit with the supervision of incorrect pixel labels of vessel. The results of DIAS show that the benefits brought by the pseudo-labeled data outweigh the drawbacks of noisy labels in IA sequence segmentation. As the number of labeled data increases, this beneficial effect gradually diminishes.
All the unlabeled images can be labeled with a new teacher, i.e., the student from the previous stage, and re-trained in the iterative training stage. Fig. 8 presents our examination of the effectiveness of iterative training in RPST. In the extremely scarce-label regime with only one labeled sequence (#1), performance is improved from 65.02% to 66.85% with the extra two-iteration training. The performance of #10 improved from 75.79% to 76.39%. Each iteration enables the model to further learn and adjust based on the foundation established by previous iterations, increasing its confidence in its own predictions and resulting in the generation of more reliable pseudo-labels, thereby enhancing segmentation accuracy. With an increasing number of training iterations, the marginal improvement in performance diminishes.
Figure 8: Result of iterative training with different amounts of labeled data.
VI Discussion
VI-A Impact of Sequences as Inputs
Figure 9: The first row is the visualization of various inputs, and the second row is the corresponding segmentation result.
The DSA sequence is a 2D temporal image sequence that represents the contrast agent’s response at different times while flowing through the blood. Theoretically, accurate delineation of vessels requires the aggregation of contrast agent responses from all images in the sequence. In clinical practice, neurosurgeons commonly utilize Minimum Intensity Projection (MIP) to process target sequences in DSA for the visualization of vascular structures [58]. MIP is a straightforward and effective method for aggregating significant vascular information from target sequences. To evaluate the impact of various input modalities on segmentation performance, we conduct comparative experiments that include subsets of the image sequence (First, Last, and First+Last), MIP and average intensity projection (AIP) image as input to the VSS-Net. As shown in Table VI, In ‘First’, where the contrast agent is exclusively displayed in the ICA, the segmentation results are inferior, the DSC is only 62.32%. In comparison, the ‘Last’ as input to the model obtained better results, thanks to the more angiographic information in thin vessel, the DSC is 76.04%. The quantitative results are consistent with those portrayed in Fig. 9 (e) and (f). When using only the first image as the model input, the large vessels are accurately segmented. However, due to the absence of angiographic information for fine vessels in the first image frame, the thin vessel at the arterial termini within the middle cerebral artery (MCA) remain incomplete. Conversely, using only the last image as input enables the accurate segmentation of most small vessels, yet it falls short in precisely delineating the larger vessels in proximity to the ICA.
TABLE VI: Results of fully-supervised segmentation with various input
Input Metrics
[…]
First image 0.6232 0.9515 0.5388 0.9864 0.4601 0.9049 12.05
Last image 0.7604 0.9647 0.7394 0.9842 0.6146 0.9813 51.89
First + Last image 0.7777 0.9668 0.7589 0.9850 0.6373 0.9846 44.22
Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 39.25
Min 0.7812 0.9673 0.7631 0.9854 0.6418 0.9842 40.02
Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
The combination of early and late arterial angiography, known as ‘First+Last’, yielded favorable results, surpassing those of both inputs used independently. As shown in Fig. 9 (a) and (b), although the combination of the ’First’ and ’Last’ images from the arterial phase can display the majority of the vascular structures filled with contrast agent, we can still utilize the intermediate images to supplement details. The experimental findings indicate that employing the complete DSA sequence as input is imperative for achieving optimal performance. The segmentation results of MIP and AIP are quite close to those obtained using the method that utilizes the full DSA sequence as input. Density projection methods such as MIP and AIP convert arterial phase sequences into 2D images along the temporal dimension, which necessitates the pre-selection of target sequences to avoid the influence of capillary and venous phase angiography. Using DSA sequences as the network’s input holds potential in addressing this challenge, as VSS-Net can selectively extract target sequence angiographic information from the entire series, including the capillary and venous phases, for the segmentation of IA.
VI-B Potential Clinical Applications and Limitations
Segmentation of vessels in DSA for quantifying their morphology has already found application in clinical research, particularly for planning interventional surgeries [7] and diagnosing strokes [5]. However, as discussed in Section VI-A, the segmentation of a single DSA frame is inadequate for fully delineating the vascular structure. We believe that the application of our proposed method for segmenting DSA vessel sequences in the aforementioned higher-level tasks will yield more accurate vascular structures. This advancement is expected to drive forward the progress of these higher-level tasks. In this work, we propose a comprehensive benchmark based on our created dataset for intracranial arterial segmentation in DSA sequences. Importantly, the proposed approach is not confined to intracranial arteries. With the availability of relevant annotated data, it can be extended to other segmentation tasks in DSA sequences, such as venous segmentation [8], as well as for the diagnosis of stroke-related conditions like aneurysms [3], ICAS, and MCAO [59].
Currently, the DIAS dataset is relatively small, primarily due to the substantial effort expended on one-to-one pixel annotations of DSA sequences. In future work, we plan to leverage RPST to generate pseudo-labels for DSA sequences pending annotation. Annotation experts will then meticulously revise these pseudo-labels to obtain precise annotations, rather than starting the annotation process from the beginning. This strategy aims to increase annotation efficiency, thereby facilitating the further expansion of DIAS. On the other hand, in the DIAS dataset, intracranial arterial sequences are selected by neurosurgeons from DSA sequences that may include stages such as the capillary or venous phase. Given the nature of 3D spatial projection, angiography from other phases in DSA may impact arterial segmentation, as the same pixel can represent an artery or a vein at different times. In our next phase of work, we plan to segment intracranial arteries from DSA sequences that include the venous phase and other stages, without relying on the neurosurgeons’ prior knowledge to pre-select arterial phase sequences.
VII Conclusion
In this paper, we present the DSA-sequence IA segmentation dataset, DIAS, which provides pixel-level annotations along with two forms of weakly supervised annotations. We establish a benchmark for evaluating IA sequence segmentation on DIAS, comprising (1) the Vessel Sequence Segmentation Network, VSS-Net; (2) the Scribble Learning-based Segmentation Framework, SSCR; and (3) the Random Patch-based Self-Training Framework, RPST. Leveraging the DIAS dataset, we have conducted comprehensive experiments to validate the effectiveness of our proposed methodologies, facilitating a deeper understanding of their performance characteristics and offering valuable insights for future research endeavors. Looking ahead, our research will persist, including the expansion of segmentation targets to venous and the segmentation of arteries from complete angiographic sequences. We emphasize that the DIAS dataset and benchmark have been made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No.62376038) and National Key R&D Program of China (No.2018AAA0102600).

### v3

@@
-DIAS: A Comprehensive Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences
+DIAS: A Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences
 Abstract
@@
 I Introduction
-Cerebrovascular diseases significantly contribute to global mortality and long-term disability [1].
-In recent decades, there has been a notable decline in the mortality rate associated with intracranial artery (IA) diseases, such as intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, largely attributed to rapid advancements in medical and imaging technology [2].
-Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3].
-It removes bony structures by subtracting the pre-contrast image, thus preserving the vessels filled with contrast agent.
-Owing to its inherently superior spatial and temporal resolution, DSA can accurately elucidate lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2].
+Cerebrovascular diseases substantially contribute to global mortality and long-term disability [1].
+Recent decades have witnessed a decline in mortality from intracranial artery (IA) diseases like intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, thanks to advancements in medical and imaging technology [2].
+Among these, Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3].
+It effectively removes bony structures by subtracting pre-contrast images, thereby enhancing the visibility of vessels filled with the contrast agent.
+Owing to its inherently superior spatial and temporal resolution, DSA can accurately reveal lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2].
 Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
-Conventionally, neuroradiologists undertake the visual analysis of DSA for diagnostic purposes, clinical decision-making, and postoperative assessment, which is inherently time-consuming, labor-intensive, and subjective.
-Automated segmentation of vessels DSA harbors the potential to mitigate this challenge by delineating vascular morphology.
+Conventionally, neuroradiologists analyze DSA images visually for diagnosis, clinical decision-making, and postoperative evaluation, which is inherently time-consuming, labor-intensive, and subjective.
+Automated vessel segmentation in DSA promises to alleviate these challenges by accurately outlining vascular morphology.
 Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7].
-Despite the undeniable significance of cerebrovascular segmentation in DSA, it remains an under-researched area.
-One possible reason is the current absence of a DSA dataset tailored specifically for the formulation and evaluation of cerebrovascular segmentation techniques in the medical imaging community.
-However, given the intrinsic attributes of its imaging modality, the task of creating a brain DSA vascular segmentation dataset is immensely challenging.
-Primarily, the DSA sequence presents a projection of intravascular contrast media from a consistent viewing angle, thereby visualizing only a fraction of the IA within the contrast agent in any given image, as illustrated in Fig.
+Despite its critical importance, cerebrovascular segmentation in DSA has not been extensively researched.
+A key factor contributing to this is the lack of a dedicated DSA dataset specifically designed for developing and evaluating segmentation techniques.
+However, the creation of such a dataset is particularly challenging due to the unique characteristics of DSA imaging.
+Primarily, in the DSA sequence, the intravascular contrast agent is projected from a fixed angle, revealing only a portion of the intracranial arteries within the contrast medium in each image, as illustrated in Fig.
 1.
-In the process of annotating vessels within DSA, neuroradiologists are required to conduct a detailed examination of each image frame, which is both laborious and expensive [8].
-Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures in imagery, diminished contrast, and multifaceted background disturbances, accentuates the difficulties associated with annotation.
+Annotating vessels in DSA necessitates neuroradiologists to meticulously examine each image frame, a process that is both labor-intensive and costly [8].
+Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast, and complex background noise, significantly increases the difficulty of accurate annotation.
 Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8].
-Yet, all these efforts are anchored to proprietary datasets, absent a consistent evaluation benchmark.
-Furthermore, most of these menthods [9, 10] are limited to single-frame image segmentation, overlooking the rich temporal characteristics inherent in DSA sequences.
-Segmentation methods based on single-frame DSA may not comprehensively capture the vascular intricacies of IA, potentially limiting their efficacy in diagnosing and treating cerebrovascular diseases [8].
-In DSA sequences, vascular segmentation is characterized as a dimensionality-reduction task, taking a 2D+time image input to yield a 2D mask output.
-This task deviates from traditional 2D and 3D vascular segmentation meathods [11, 12, 13].
-The pivotal challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
-Figure 1: A representative sample of intracranial vessels in sagittal view from DSA, where t to t+5 are consecutive frames of the arterial phase.
-In this work, we construct the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons.
-Given the intricate and labor-intensive nature of vessel annotation within DSA sequences, we’ve provided two types of weak annotations for training samples, aiming to facilitate studies on weakly-supervised IA segmentation.
-For the DIAS, we have devised three benchmarks tailored for fully-supervised segmentation (FSS), weakly-supervised segmentation (WSS), and semi-supervised segmentation (SSS): the vessel sequence segmentation network, the scribble learning-based segmentation framework and the random patch-based self-training framework.
-We have carried out comprehensive experiments using the DIAS, establishing robust baselines that will serve as a foundation for future algorithmic comparisons on the DIAS dataset.
-We hope that the DIAS dataset could facilitate open research in cerebral vessel segmentation.
+However, these studies rely on in-house datasets, lacking a unified standard for evaluation.
+Most of these menthods [9, 10] focus on segmenting single-frame images, thereby ignoring DSA sequences’ inherent temporal attributes.
+As a result, single-frame based segmentation approaches may not fully represent the intricate vascular morphology of IA, potentially compromising their effectiveness in the diagnosis and treatment of cerebrovascular diseases [8].
+Vessel segmentation in DSA is considered a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map.
+This approach deviates from the traditional 2D and 3D vessel segmentation methods [11, 12, 13].
+The key challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
+Figure 1: A representative sample of intracranial vessels in lateral view from DSA, where t to t+5 are consecutive frames of the arterial phase.
+The red arrow in the upper left image indicates segments C1-C3 of the intracranial internal carotid artery (ICA), while the lower left image represents the anterior cerebral artery (ACA) and the middle cerebral artery (MCA).
+In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons.
+Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation.
+To comprehensively evaluate the efficacy of prevalent segmentation methods like fully-supervised, weakly-supervised, and semi-supervised segmentation on the DIAS dataset, we have established a benchmark comprising a vessel sequence segmentation network, a scribble-based learning segmentation framework, and a random patch-based self-training semi-supervised segmentation pipeline.
+Extensive experiments with the DIAS set robust baselines, potentially advancing future algorithmic developments in cerebral vessel segmentation.
 TABLE I: Summary of several public vessel segmentation datasets. ${}^{\dagger}$ represents only training set data is available.
@@
 II-A Vessel Segmentation Datasets
-Vessel segmentation is a popular area of research within computer-aided medicine, which has witnessed remarkable progress in recent years, owing to technological advancements and the availability of high-quality annotated datasets.
+Vessel segmentation, a prominent research field in computer-aided medicine, has seen significant advancements recently, largely due to technological progress and the availability of annotated datasets.
 We have summarized the several publicly available vessel segmentation datasets in Table I.
-A representative example is retinal vessel segmentation, which has many publicly available datasets due to its easy availability, such as colour fundus datasets (DRIVE [14], CHASE_DB1 [15], STARE [16], HRF [17]) and OCTA datasets (ROSE1 and ROSE2 [18]).
+A representative example is retinal vessel segmentation, which has many publicly available datasets, such as colour fundus datasets (DRIVE [14], CHASE_DB1 [15], STARE [16], and HRF [17]) and OCTA datasets (ROSE1 and ROSE2 [18]).
 These datasets consist of 20-117 2D images with spatial resolution ranging from 304$\times$304 to 3504$\times$2336.
@@
 ASOCA [21] and VESSEL12 [22] are the 3D vessel segmentation datasets of coronary artery CTA and lung CT from the challenges, in which only training set data is available.
-The availability of such datasets has led to a surge of outstanding works in the field of segmentation, with researchers striving to develop algorithms that outperform existing methods.
-So far, there have been no publicly available DSA datasets with vessel annotation in the medical imaging community, and only a few studies conducted with in-house datasets.
+The proliferation of such datasets has catalyzed a wave of innovative research in the area of segmentation, with scholars endeavoring to devise algorithms that surpass current methodologies.
+Until now, the medical imaging community lacked publicly available DSA datasets with vessel annotations, with only limited studies using proprietary data.
 DCVessel [9] comprises 30 DSA arterial-phase images at a resolution of 700$\times$700, while HEART [23] comprises 1092 DSA coronary images at 256$\times$256 resolution.
 Su et al.
-[8] manually segmented arteries and veins on 97 DSA series, covering both anterior-posterior and lateral views, each with a resolution of 1024$\times$1024 pixels.
-In this work, we have released DIAS, an IA segmentation dataset derived from DSA sequences.
-It comprises 120 sequences (753 frames), each with a resolution of 800$\times$800 pixels, and the lengths of these sequences range from 4 to 14.
-The DIAS can serve as a reliable benchmark for the development of algorithms on IA segmentation from DSA sequences.
-We hope that this dataset will intrigue researchers to explore this field further.
+[8] manually segmented arteries and veins on 97 DSA series, covering both anteroposterior and lateral views, each with a resolution of 1024$\times$1024 pixels.
+Our contribution, DIAS, offers a comprehensive IA segmentation dataset from DSA sequences, comprising 120 sequences (753 frames) at 800$\times$800 pixel resolution, spanning lengths of 4 to 14 frames.
+The DIAS can serve as a reliable benchmark for the development of algorithms on IA segmentation, hopefully inspiring further research in this arena.
 II-B Vessel Segmentation Methods
-In the past few years, deep learning has demonstrated great potential in vessel segmentation, surpassing traditional methods, such as hand-crafted features [24], filtering-based models [25], and statistical models [26].
-Deep learning-based methods have become primary approaches for vessel segmentation, with UNet [27] serving a crucial role as a milestone network architecture, giving rise to numerous variant networks.
-These networks utilize an encoder-decoder architecture with popular components, such as multi-scale design [13, 9, 28], deep supervision [29, 9, 12], attention mechanisms [30, 13, 31], atrous convolution [32, 29, 12, 9], and improved skip connections [9, 12].
+In the past few years, deep learning has revolutionized vessel segmentation, outperforming traditional approaches like hand-crafted features [24], filtering-based models [25], and statistical models [26].
+Deep learning-based methods, particularly UNet [27] and its numerous variants, have become the primary methods for vessel segmentation.
+These networks typically employ an encoder-decoder architecture with popular components, such as multi-scale design [13, 9, 28], deep supervision [29, 9, 12], attention mechanisms [30, 13, 31], atrous convolution [32, 29, 12, 9], and improved skip connections [9, 12].
 Moreover, several novel neural networks [33, 34] have also succeeded in vessel segmentation by leveraging generative adversarial networks (GAN) or graph neural networks (GNN).
-Some studies focused on vessel segmentation beyond network architecture, such as optimized training strategies [35, 36] and modifying loss functions [37].
+Beyond network architectures, some research has focused on optimizing training strategies [35, 36] and modifying loss functions [37].
 However, there have been few studies on DSA vessel segmentation.
 Meng et al.
-[9] proposed a CNN-based multiscale dense CNN segmentation framework, which includes a multiscale atrous convolution module, an improved dense block, and redesigned skip connections, for IA segmentation in single frame DSA.
+[9] proposed a CNN-based multiscale dense CNN segmentation network for IA segmentation in single-frame DSA, utilizing multiscale atrous convolution, enhanced dense blocks, and redesigned skip connections.
 Zhang et al.
@@
 Su et al.
-proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8].
+proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8].
 While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
 Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers.
-WSL trains models with sparse annotations instead of full ones.
+Weakly-supervised segmentation (WSS) trains models with sparse annotations instead of full ones.
 The common methods involves online labeling [39] via consistency checks with the mean teacher or using pseudo-labels supervision [40] derived from model predictions.
-SSS aims to explore tremendous unlabeled data with supervision from limited labeled data.
+Semi-supervised segmentation (SSS) aims to explore tremendous unlabeled data with supervision from limited labeled data.
 Recently, self-training methods [41, 42] have dominated this field.
@@
 To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences.
-In this regard, we propose a comprehensive benchmark involving FSS, WSS, and SSS for IA segmentation based on the constructed DSA-sequences dataset.
-Figure 2: Full annotated process of DIAS.
-Figure 3: Full annotated representative samples of DIAS at coronal view and sagittal view.
-Figure 4: Examples of two forms scribble annotations: the scribbles annotation with little clinical experience (SALE) and Random Drawing based the Full Annotation (RDFA).
-Red represents annotated vessel pixels, blue represents annotated background pixels.
+Building upon our created dataset, we have conducted in-depth research and proposed innovative methods in three distinct areas: network models, WSS, and SSS.
+These approaches are specifically tailored to leverage the unique features of the dataset, addressing each segmentation paradigm comprehensively.
 III DIAS Dataset
 III-A Dataset Description
-The DIAS dataset was derived from intraoperative DSA images obtained during ischemic stroke treatments, including ICAS and MCAO, at Beijing Tiantan Hospital.
+The DIAS dataset is derived from DSA images obtained during ischemic stroke treatments at Beijing Tiantan Hospital.
+This dataset encompasses cases of ICAS and MCAO.
 The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study.
 All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details.
-The original DSA sequences were captured by the angiography machine during the actual endovascular therapy process and stored as DICOM files.
-It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view, lateral view or both.
-In this study, we retained 120 sequences after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates.
+The original DSA sequences were captured by the Siemens angiography machine during the neurointerventional procedures and stored as DICOM files.
+It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view or lateral view.
+In this study, we retained 120 sequences (753 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates.
+A total of 120 sequences were derived from 60 patients, among whom 24 were female and 36 were male.
+The average age of these patients was 62.5 years, with an age range from 40 to 88 years old.
 The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS).
@@
 III-B1 Full Annotation
-Figure 5: Illustration of the proposed pipeline for DSA-sequence intracranial artery segmentation.
-(a) Vessel sequence segmentation network with (b) sequence feature extraction module; (c) WSL: scribble learning-based segmentation composed of scribble supervision and consistency regularization; (d) SSL: random patch-based self-training framework.
-(WSC: Weight Sharing Convolution, SF: Sequence Fusion, GMP: Global Max Pooling, WSD: Weight Sharing Downsampling, DA: Data augmentation, WA: Weak Augmentation, SA: Strong Augmentation)
-Manual annotation of IA in DSA sequences requires taking into account all the image information in the sequence.
+Figure 2: Full annotated process of DIAS.
+The manual annotation of IA in DSA sequences necessitates a thorough consideration of all image data within the sequence.
 Fig.
-1 shows an example of intracranial vessels in five consecutive frames of the arterial phase in the sagittal view.
-During this period, the contrast agent is gradually dispersed from the internal carotid artery (ICA) to cerebral vessels such as the anterior cerebral artery (ACA), middle cerebral artery (MCA), and posterior cerebral artery (PCA) with blood flow.
-For a single image, the contrast agent is visible only in a portion of the IA.
-For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+4, it filled the ACA, MCA, and PCA.
-Annotating the IA of each image in the sequence and then merging the annotation mask can maximize the utilization of the sequence’s vessel information to obtain accurate annotations.
-The thinness of vessels, particularly in later-stage arterial images, renders the commonly used strategy of outlining their borders in medical image annotation unfeasible.
+1 illustrates a typical representation of intracranial vessels across five successive frames during the arterial phase in a lateral view.
+During this phase, the contrast agent progressively spreads from the internal carotid artery (ICA) into cerebral vessels like the anterior cerebral artery (ACA) middle cerebral artery (MCA), carried by the blood flow.
+For a single image, the contrast agent is discernible in only a portion of the IA.
+For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+3, it filled the ACA and MCA.
+To annotate the IA effectively, each image in the sequence must be individually annotated, and then these annotations should be integrated to leverage the full vessel information available in the sequence, thereby ensuring precision in annotation.
+However, the challenge arises with the fine nature of vessels, especially in images from later arterial stages, which makes the standard approach of demarcating their borders impractical for medical image annotation.
 Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure.
+In reality, the angiographic representations of vessels in adjacent frame images exhibit a high degree of similarity, indicating that annotating every pixel in each frame image involves a substantial amount of repetitive work.
 Fig.
-3 shows that our proposal for a compromise solution, which reduces workload while maintaining annotation accuracy:
-  • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the later phase.
-The contrast agent during the anterior phase predominantly fills the ICA as opposed to the smaller vessels such as the ACA, MCA, and PCA arteries, which are filled by the contrast agent in the later phase.
+2 presents our proposed solution, a pragmatic compromise that aims to reduce the workload without compromising the accuracy of annotations:
+Figure 3: Full annotated representative samples of DIAS at anteroposterior view and lateral view.
+Figure 4: Examples of two forms scribble annotations: the Scribbles Annotation with Little clinical Experience (SALE) and Random Drawing based the Full Annotation (RDFA).
+Red represents annotated vessel pixels, blue represents annotated background pixels.
+  • [leftmargin=*,itemsep=-4pt]
+  • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the posterior phase.
+The contrast agent during the anterior phase predominantly fills the ICA as opposed to the smaller vessels such as the ACA and MCA arteries, which are filled by the contrast agent in the posterior phase.
   • The last image of each phase is annotated pixel-wise, followed by a meticulous review of other same-phase images to finalize the annotation.
 Thick vessels are the focus of annotation in the anterior phase, while thin vessels are prioritized in the later phase.
-  • The annotations of thick and thin vessels are merged using a pixel-level OR operation (1 for vessel pixel and 0 for background) to create the ultimate vessel mask.
-We randomly selected 60 sequences from DIAS for the purpose of annotating blood vessels, using the patient ID as the criterion for selection.
-The remaining 60 sequences were used for research on IA semi-supervised segmentation.
-The DSA sequence was initially annotated by two MD students of surgery, followed by corrections from an associate physician with 10 years of experience, and finally confirmed and double-checked by a chief physician with 15 years of experience.
+  • The annotations of thick and thin vessels are merged using a pixel-level OR operation (1 for vessel pixel and 0 for background) to create the ultimate groudtruth.
+We randomly selected 60 sequences from the DIAS dataset for the purpose of annotating vessels, using patient IDs as the criterion for selection.
+The remaining 60 sequences were used for research on semi-supervised segmentation of IA.
+The DSA sequence was initially annotated by two MD students specializing in neurosurgery, followed by corrections from an associate physician with 10 years of experience, and finally confirmed and double-checked by a chief physician with 15 years of experience.
+Fig.
+3 presents fully annotated examples of intracranial arteries in anteroposterior and lateral views.
+Figure 5: Illustration of the proposed benchmark for DSA-sequence intracranial artery segmentation.
+(a) Vessel sequence segmentation network with (b) sequence feature extraction module; (c) WSS: scribble learning-based segmentation composed of scribble supervision and consistency regularization; (d) SSS: random patch-based self-training framework.
+(WSC: Weight Sharing Convolution, SF: Sequence Fusion, GMP: Global Max Pooling, WSD: Weight Sharing Downsampling, DA: Data augmentation, WA: Weak Augmentation, SA: Strong Augmentation)
 III-B2 Scribble Annotation
@@
 This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries.
-To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based the Full Annotation (RDFA).
+To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based on the Full Annotation (RDFA).
 SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig.
@@
 Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging.
-Consequently, pixels in these regions were annotated during annotation.
+Consequently, pixels in these regions were not annotated during annotation.
 Moreover, scribble segmentation relies on supervised information for every category.
@@
 Automatic IA segmentation in DSA sequences is an under-explored and challenging task.
-To fully explore potential research topics, three elaborately designed benchmarks for IA segmentation were proposed, aiming to provide comprehensive benchmarks for FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
+To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
 IV-A Fully-supervised Segmentation
 Segmenting IA in DSA sequence can be considered as a dimensionality reduction (DR) task.
-Given an image $I\in\mathbb{R}^{\prod_{d=1}^{N}n_{d}}$, where $N$ represents the number of dimensions and $n_{d}$ is the size of the image in the corresponding dimension $d$, DR segmentation can be defined as finding a function $f$ such that $f:I\rightarrow O$, where $O\in\mathbb{R}^{\prod_{d=1}^{M}n_{d}}$ represents segmentation output and $M$ ($<N$) is its dimensionality [47].
-For IA segmentation in DSA sequence, $N=3$, $M=2$.
-Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representation of vessels from various scales to achieve precise IA segmentation in DSA.
+Given an image $I\in\mathbb{R}^{\prod_{d=1}^{M}n_{d}}$, where $M$ represents the number of dimensions and $n_{d}$ is the size of the image in the corresponding dimension $d$, DR segmentation can be defined as finding a function $f$ such that $f:I\rightarrow O$, where $O\in\mathbb{R}^{\prod_{d=1}^{N}n_{d}}$ represents segmentation output and $N$ ($<M$) is its dimensionality [47].
+For IA segmentation in DSA sequence, $M=3$, $N=2$.
+Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA.
 Figure 5 (a,b) illustrates an overview of the proposed network architecture.
-VSS-Net is composed of multi-scale sequence feature extraction modules (SFEM).
-These modules handle the 2D+T input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations.
+VSS-Net comprises multi-scale sequence feature extraction modules (SFEM).
+These modules handle the 2D+Time input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations.
 Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
@@
 Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation.
-Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1x1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels.
+Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels.
 This process is formulated as:
-$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i})$ (1)
-where, $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence.
+$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i}),$ (1)
+where $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence.
 $W$ are $1\times 1$ convolutional kernels.
@@
 $\oplus$ denotes the concatenation operation.
-Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of sequence.
+Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of the sequence.
 ConvGRU [48] is an extension of the traditional GRU that integrates convolutional operations, making it more suitable for spatial data like images.
 Consider $x_{t}$ as the input to the ConvGRU, with corresponding hidden states denoted by $h_{t}$, we can delineate the updating process of ConvGRU unrolled by time as follows:
-$z_{t}$ $=\sigma(W_{z}\ast x_{z}+U_{z}\ast h_{t-1})$ (2)
-$r_{t}$ $=\sigma(W_{r}\ast x_{r}+U_{r}\ast h_{t-1})$ (3)
-$\tilde{h}_{t}$ $=\tanh(W_{h}\ast x_{t}+U_{h}\ast(r_{t}\odot h_{t-1}))$ (4)
-$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (5)
-where, $\sigma$ represents the sigmoid function.
+$z_{t}$ $=\sigma(W_{z}\ast x_{z}+U_{z}\ast h_{t-1}),$ (2)
+$r_{t}$ $=\sigma(W_{r}\ast x_{r}+U_{r}\ast h_{t-1}),$ (3)
+$\tilde{h}_{t}$ $=\tanh(W_{h}\ast x_{t}+U_{h}\ast(r_{t}\odot h_{t-1})),$ (4)
+$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t},$ (5)
+where $\sigma$ represents the sigmoid function.
 $\tanh$ represents the hyperbolic tangent function.
@@
 The process can be formulated as follows:
-$h_{t}^{f}$ $=ConvGRU(h_{t-1},x_{t})$ (6)
-$h_{t}^{b}$ $=ConvGRU(h_{t+1},x_{t})$ (7)
-$h_{t}$ $=\tanh(W_{f}\ast h_{t}^{f}+W_{b}\ast h_{t}^{b})$ (8)
-where, $h_{t}^{f}$ and $h_{t}^{b}$ denote the hidden states from the forward and backward ConvGRU units, respectively. $x_{t}$ is the $t^{th}$ output feature from WSC.
+$h_{t}^{f}$ $=ConvGRU(h_{t-1},x_{t}),$ (6)
+$h_{t}^{b}$ $=ConvGRU(h_{t+1},x_{t}),$ (7)
+$h_{t}$ $=\tanh(W_{f}\ast h_{t}^{f}+W_{b}\ast h_{t}^{b}),$ (8)
+where $h_{t}^{f}$ and $h_{t}^{b}$ denote the hidden states from the forward and backward ConvGRU units, respectively. $x_{t}$ is the $t^{th}$ output feature from WSC.
 The output feature maps are spatially downsampled using stride convolution with weight sharing to serve as inputs for the next stage encoder.
@@
 IV-B Weakly-supervised Segmentation
-A common strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled points.
-However, this baseline approach does not utilize the unlabeled points that contain critical semantic information.
-The performance degradation is unavoidable due to the lack of labeled data points.
+A prevalent strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled pixels.
+However, this baseline approach fails to harness the unlabeled points that contain critical semantic information.
+Consequently, the absence of sufficient labeled data invariably results in performance degradation.
 To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels.
@@
 Fig.
-5 (b) presents the implementation framework for SSCR.
+5 (c) presents the implementation framework for SSCR.
 SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
@@
 The function, namely partial cross entropy (pCE), can be formulated as:
-$\mathcal{L}_{pce}(y,\hat{y})=-\sum_{i=1}^{M}\sum_{j=1}^{N}{\mathbbm{1}(p_{i,j}\in S)(y_{i,j}\log(\hat{y}_{i,j|\theta})})$ (9)
-where, $M$ and $N$ represent the width and height of the image, respectively.
+$\mathcal{L}_{pce}(y,\hat{y})=-\sum_{i=1}^{M}\sum_{j=1}^{N}{\mathbbm{1}(p_{i,j}\in S)(y_{i,j}\log(\hat{y}_{i,j|\theta})}),$ (9)
+where $M$ and $N$ represent the width and height of the image, respectively.
 $\hat{y}_{i,j|\theta}$ is the predict for the pixel $p_{i,j}\in S$ given the network parameters $\theta$, and $y_{i,j}$ denotes the ground truth label.
@@
 $PL_{i}=\begin{cases}s_{i,j},&\text{if}\ p_{i,j}\in S\\ \text{argmax}(\text{softmax}(\mathcal{H}(p_{i,j};\theta)[C])),&\text{otherwise}\par\end{cases}$ (10)
-where, $s_{i,j}\in S$ defines annotated pixels, and $\mathcal{H}(p_{i,j};\theta)$ defines the predicted output of $p_{i,j}$. $C$ is the channel of annotated class, i.e., 0 and 1.
+where $s_{i,j}\in S$ defines annotated pixels, and $\mathcal{H}(p_{i,j};\theta)$ defines the predicted output of $p_{i,j}$. $C$ is the channel of annotated class, i.e., 0 and 1.
 We enforce network consistency by utilizing cross pseudo supervision, which employs pseudo labels from one network to supervise the other network [43].
@@
 The training objective contains three losses: scribble supervision loss $\mathcal{L}_{scr}$, cross pseudo supervision loss $\mathcal{L}_{cps}$, and consistency loss $\mathcal{L}_{con}$:
-$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL))$ (11)
-$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A}))$ (12)
-$\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B})$ (13)
-where, $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function.
+$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL)),$ (11)
+$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A})),$ (12)
+$\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B}),$ (13)
+where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function.
 $P_{A}$ and $P_{B}$ are the outputs of two models.
@@
 $SL$ is the scribble label.
-Finally, the training objective $\mathcal{L}$ is formulated as
-$\mathcal{L}=\mathcal{L}_{scr}+\lambda_{1}\mathcal{L}_{cps}+\lambda_{2}\mathcal{L}_{con}$ (14)
-where, $\lambda_{1}$ and $\lambda_{2}$ are the weight of loss components $\mathcal{L}_{cps}$ and $\mathcal{L}_{con}$.
+Finally, the training objective $\mathcal{L}$ is formulated as:
+$\mathcal{L}=\mathcal{L}_{scr}+\lambda_{1}\mathcal{L}_{cps}+\lambda_{2}\mathcal{L}_{con},$ (14)
+where $\lambda_{1}$ and $\lambda_{2}$ are the weight of loss components $\mathcal{L}_{cps}$ and $\mathcal{L}_{con}$.
 IV-C Semi-supervised Segmentation
@@
 The primary challenge in semi-supervised setting lies in how to effectively utilize the unlabeled images.
-Self-training [50] via pseudo labeling is a conventional, simple, and popular pipeline to leverage unlabeled data, where the retrained student is supervised with hard labels produced by the teacher trained on labeled data, which is commonly regarded as a form of entropy minimization in semi-supervised learning.
+Self-training [50] via pseudo labeling is a conventional, simple, and popular pipeline to leverage unlabeled data, where the retrained student is supervised with hard labels produced by the teacher trained on labeled data.
 For semi-supervised segmentation problem, let $\mathcal{D}^{l}=\{(x_{i},y_{i})\}_{i=1}^{M}$ be pixel-wise labeled images and $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$ be unlabeled images, where in most cases $N\gg M$.
 Self-training can be simplified into the following steps:
-  1. [leftmargin=*,itemsep=-5pt]
+  1. [leftmargin=*,itemsep=-4pt]
   2. 1.
@@
     Re-train a student model $S$ on the label data and pseudo-label data $\mathcal{D}^{l}\cup\mathcal{\hat{D}}^{u}$ for final evaluation.
-We follow the plainest self-training strategy to provide a SSS benchmark for DIAS.
+As shown in Fig.
+5 (d), we follow the plainest self-training strategy to provide a SSS benchmark for DIAS.
 The injection of strong data augmentations on unlabeled images is highly advantageous as it decouples the predictions as well as alleviates overfitting on noisy pseudo-labels.
 Despite its simplicity, self-training with strong data augmentations surpasses existing methods without any bells and whistles [41].
-We performed two types of data augmentations in our implementation: weak and strong, denoted by $A^{s}$ and $A^{w}$ respectively.
+We performed two types of data augmentations in our implementation: weak and strong, denoted by $A^{w}$ and $A^{s}$ respectively.
 Weak data augmentations (WDA) consist of horizontal flipping, vertical flipping, and $[90,180,270]$ degree rotation.
 Strong data augmentations (SDA) include Cutout, adding white Gaussian noise, Gaussian blurring, and elastic deformation.
-We randomly adjust the contrast of each image in the DSA sequence to simulate angiography with intensity differences.
+Furthermore, we propose the application of random contrast adjustments to each image within the sequence to simulate angiography with intensity differences.
 We randomly extract patches from the DSA sequence to train the teacher and student models.
@@
           Unlabeled training set $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$,
-          Weak/strong augmentation $\mathcal{A}^{w}/{A}^{s}$,
+          Weak/strong augmentation $\mathcal{A}^{w}/\mathcal{A}^{s}$,
           Teacher/student model $T/S$,
@@
 V-A Implementation Details and Evaluation Metrics
+TABLE II: Results of fully-supervised segmentation with state-of-the-art models.
+Methods Params / M Metrics
+[…]
+2D UNet [27] 31.04 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
+Res-UNet [51] 9.37 0.7656 0.9655 0.7425 0.9850 0.6235 0.9816 60.56
+UNet++ [52] 9.05 0.7768 0.9659 0.7726 0.9830 0.6374 0.9830 54.65
+Att-UNet [53] 8.73 0.7760 0.9660 0.7686 0.9835 0.6258 0.9811 59.37
+[…]
+MAA-Net [13] 5.80 0.7693 0.9661 0.7423 0.9855 0.6361 0.9818 68.18
+FR-UNet [12] 7.37 0.7779 0.9665 0.7646 0.9841 0.6376 0.9827 53.92
+3D UNet [27] 90.27 0.7691 0.9648 0.7690 0.9824 0.6275 0.9819 62.92
+Res-UNet [51] 26.38 0.7701 0.9653 0.7597 0.9835 0.6284 0.9817 73.05
+UNet++ [52] 26.17 0.7649 0.9639 0.7711 0.9811 0.6214 0.9811 67.02
+Att-UNet [53] 23.6 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
+[…]
+FR-UNet [12] 20.57 0.7635 0.9646 0.7518 0.9830 0.6210 0.9815 65.67
+[…]
+PSC [47] 6.43 0.7649 0.9646 0.7552 0.9832 0.6223 0.9814 76.04
+SVS-Net [38] 8.48 0.7698 0.9659 0.7474 0.9853 0.6277 0.9803 97.28
+ST-UNet [8] 104.37 0.7818 0.9675 0.7533 0.9864 0.6436 0.9843 62.59
+VSS-Net 18.69 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
 The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level.
 To facilitate input into the model, all DSA sequences were resampled to a length of 8.
-All DSA sequences were z-score normalized.
+Furthermore, z-score normalization was applied across all DSA sequences for standardization.
 Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64.
@@
 The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs.
-The parameter $P$ was set to 0.5, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment.
-For SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5 .
+The parameter $P$ was set to 0.5 in the RPST, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment.
+For the SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5.
 We only use one network to generate the results for evaluation.
@@
 […]
-$y=\mathcal{H}_{2d}(Squ(x))$ (15)
-where, $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
+$y=\mathcal{H}_{2d}(Squ(x)),$ (15)
+where $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
 For a 3D model $\mathcal{H}_{3d}$, the DSA sequence is fed directly into the model, where it is treated as volume data.
@@
 Finally, $1\times 1$ convolution reduces $F$ to the number of classes:
-$y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x)))$ (16)
+$y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x))).$ (16)
 Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods.
@@
 PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures.
+Figure 6: Visualization of segmentation results of state-of-the-art models.
+The first and third rows display images of the anteroposterior view, while the second and fourth rows present images of the lateral view.
+In the segmentation maps for various methods, red pixels indicate false negatives, and green pixels signify false positives.
+The white and black pixels represent correctly segmented vessels and background, respectively.
 Upon overall comparison of the DSC scores for all ‘DR+2D model’ and ‘DR+3D model’, 2D models slightly outperform 3D models with fewer parameters.
@@
 However, the approach of setting fixed channel numbers is not applicable to variable image sequences.
-We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Figure 6.
+We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Fig.
+6.
 In the figure, the red and blue colors represent pixels that are false negatives and false positives, respectively, indicative of segmentation errors.
 It can be observed that a prevalent issue shared by all methods is the abundant incorrect segmentation of thin vessels.
-Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Figure 6.
+Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Fig.
+6.
 In comparison with other methods, VSS-Net has the highest number of white pixels in the fine vessels, indicating the best segmentation results.
-TABLE II: Results of fully-supervised segmentation with state-of-the-art models.
-Methods Params / M Metrics
-[…]
-2D UNet [27] 31.04 0.7677 0.9647 0.757 0.9829 0.6258 0.9811 68.29
-Res-UNet [54] 9.37 0.7656 0.9655 0.7425 0.985 0.6235 0.9816 60.56
-UNet++ [53] 9.05 0.7768 0.9659 0.7726 0.983 0.6374 0.983 54.65
-Att-UNet [52] 8.73 0.7760 0.9660 0.7686 0.9835 0.6258 0.9811 59.37
-[…]
-MAA-Net [13] 5.80 0.7693 0.9661 0.7423 0.9855 0.6361 0.9818 68.18
-FR-UNet [12] 7.37 0.7779 0.9665 0.7646 0.9841 0.6376 0.9827 53.92
-3D UNet [27] 90.27 0.7691 0.9648 0.769 0.9824 0.6275 0.9819 62.92
-Res-UNet [54] 26.38 0.7701 0.9653 0.7597 0.9835 0.6284 0.9817 73.05
-UNet++ [53] 26.17 0.7649 0.9639 0.7711 0.9811 0.6214 0.9811 67.02
-Att-UNet [52] 23.6 0.7677 0.9647 0.757 0.9829 0.6258 0.9811 68.29
-[…]
-FR-UNet [12] 20.57 0.7635 0.9646 0.7518 0.983 0.621 0.9815 65.67
-[…]
-PSC [47] 6.43 0.7649 0.9646 0.7552 0.9832 0.6223 0.9814 76.04
-SVS-Net [38] 8.48 0.7698 0.9659 0.7474 0.9853 0.6277 0.9803 97.28
-ST-UNet [8] 104.37 0.7818 0.9675 0.7533 0.9864 0.6436 0.9843 62.59
-VSS-Net 18.69 0.7822 0.9668 0.777 0.9838 0.6442 0.9857 43.02
-Figure 6: Visualization of segmentation results of state-of-the-art models.
-The first row is coronal view, and the second row is sagittal view.
-The red pixel indicates false negatives, and the green pixel indicates false positives.
 TABLE III: Results of Scribble-supervised IA sequence segmentation.
@@
 V-C Weakly-supervised Segmentation
-We compared our method with five scribble-supervised segmentation methods: partial Cross Entropy (pCE) [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40].
+We compared our method with five scribble-supervised segmentation methods: pCE [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40].
 To ensure a fair comparison, all experimental networks employed the UNet [27], consistent with the approach in DMPLS.
@@
 Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 107.35
-Student w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 108.96
+S w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 108.96
 Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 103.09
-Teacher 3 0 0.6919 0.949 0.7358 0.9673 0.5336 0.9182 320.02
+Teacher 3 0 0.6919 0.9490 0.7358 0.9673 0.5336 0.9182 320.02
 Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 121.84
-Student w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.57 0.9578 125.83
+S w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.5700 0.9578 125.83
 Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 125.78
-Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.591 0.9694 148.89
-Student 10 30 0.7544 0.9618 0.766 0.9786 0.6087 0.9769 93.31
-Student w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
-Student 10 60 0.7579 0.9625 0.7662 0.9795 0.613 0.9786 87.72
-We conducted ablation studies to investigate the impact of various components in our proposed pipeline for weakly-supervised IA segmentation.
+Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 148.89
+Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 93.31
+S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
+Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 87.72
+We conducted ablation studies to investigate the impact of various components in our proposed framework for weakly-supervised IA segmentation.
 As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis.
@@
 This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability.
-We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation without CPS.
+We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation.
 Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC.
@@
 Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 39.25
-Min 0.7812 0.9673 0.7631 0.9854 0.6418 0.9842 40.02
+Min 0.7802 0.9673 0.7631 0.9854 0.6418 0.9842 40.02
 Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
@@
 However, as discussed in Section VI-A, the segmentation of a single DSA frame is inadequate for fully delineating the vascular structure.
-We believe that the application of our proposed method for segmenting DSA vessel sequences in the aforementioned higher-level tasks will yield more accurate vascular structures.
+We believe that the application of VSS-Net for segmenting DSA vessel sequences in the aforementioned higher-level tasks will yield more accurate vascular structures.
 This advancement is expected to drive forward the progress of these higher-level tasks.
-In this work, we propose a comprehensive benchmark based on our created dataset for intracranial arterial segmentation in DSA sequences.
+Although the SSCR using scribble annotations falls short compared to fully supervised segmentation, the results shown in Fig.
+7 reveal that this shortfall is primarily in the segmentation of smaller, terminal branches of vessels.
+Conversely, the segmentation of main vessels such as ICA, ACA, and MCA is comparatively effective.
+This is of clinical importance, as stroke-related ailments such as aneurysms [3], ICAS, and MCAO [59] predominantly occur in these main vessels.
+It is noteworthy that our assessment indicates the time taken for scribble annotation of a DSA sequence is just one-tenth of that required for full annotation.
+In this work, we introduce the FSS, WSS, and SSS methods for intracranial artery segmentation in DSA sequences, specifically designed for the DIAS dataset.
 Importantly, the proposed approach is not confined to intracranial arteries.
-With the availability of relevant annotated data, it can be extended to other segmentation tasks in DSA sequences, such as venous segmentation [8], as well as for the diagnosis of stroke-related conditions like aneurysms [3], ICAS, and MCAO [59].
+With the availability of relevant annotated data, it can be extended to other segmentation tasks in DSA sequences, such as venous segmentation [8], as well as for the diagnosis of stroke-related conditions like aneurysms, ICAS, and MCAO.
+The experimental results of the RPST demonstrates the potential of semi-supervised segmentation methods in the segmentation of IA in DSA sequences.
+One straightforward application is in assisting with data annotation.
 Currently, the DIAS dataset is relatively small, primarily due to the substantial effort expended on one-to-one pixel annotations of DSA sequences.
@@
 We emphasize that the DIAS dataset and benchmark have been made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.
-Acknowledgments
-This work was supported in part by the National Natural Science Foundation of China (No.62376038) and National Key R&D Program of China (No.2018AAA0102600).

### v4

@@
 Abstract
-Digital subtraction angiography (DSA) is universally acknowledged as the gold standard for examining lesion angioarchitecture, elucidating arterial blood supply dynamics, and guiding endovascular interventions.
-The automatic segmentation of intracranial arteries (IA) in DSA, which is pivotal for quantifying vascular morphology, plays an essential role in computer-assisted stroke research and clinical practices.
-Nevertheless, research in this specific domain remains constrained, primarily owing to the unavailability of publicly datasets for IA segmentation within the research community.
-Currently, the predominant focus of methodologies lies in the segmentation of single-frame DSA using in-house datasets.
-These methods, limited by the partial inclusion of contrast in single-frame DSA, encounters challenges in rendering a precise representation of vascular structures.
-In this paper, we introduces DIAS, a dataset specifically developed for IA segmentation in DSA sequences.
-A comprehensive benchmark has been established for evaluating DIAS, covering fully, weakly, and semi-supervised segmentation methods.
-Specifically, we propose a vessel sequence segmentation network that captures the spatiotemporal representation of intravascular contrast for segmenting vessels in DSA sequences.
-For weakly-supervised learning, we propose a novel scribble learning-based image segmentation framework, incorporating both scribble supervision and consistency regularization.
-Furthermore, we introduce a random patch-based self-training framework that harnesses unlabeled DSA sequences to improve segmentation performance.
+The automated segmentation of Intracranial Arteries (IA) in Digital Subtraction Angiography (DSA) plays a crucial role in the quantification of vascular morphology, significantly contributing to computer-assisted stroke research and clinical practice.
+Current research primarily focuses on the segmentation of single-frame DSA using proprietary datasets.
+However, these methods face challenges due to the inherent limitation of single-frame DSA, which only partially displays vascular contrast, thereby hindering accurate vascular structure representation.
+In this work, we introduce DIAS, a dataset specifically developed for IA segmentation in DSA sequences.
+We establish a comprehensive benchmark for evaluating DIAS, covering full, weak, and semi-supervised segmentation methods.
+Specifically, we propose the vessel sequence segmentation network, in which the sequence feature extraction module effectively captures spatiotemporal representations of intravascular contrast, achieving intracranial artery segmentation in 2D+Time DSA sequences.
+For weakly-supervised IA segmentation, we propose a novel scribble learning-based image segmentation framework, which, under the guidance of scribble labels, employs cross pseudo-supervision and consistency regularization to improve the performance of the segmentation network.
+Furthermore, we introduce the random patch-based self-training framework, aimed at alleviating the performance constraints encountered in IA segmentation due to the limited availability of annotated DSA data.
 Our extensive experiments on the DIAS dataset demonstrate the effectiveness of these methods as potential baselines for future research and clinical applications.
-The dataset and code will be publicly available at https://github.com/lseventeen/DIAS.
+The dataset and code are publicly available at https://doi.org/10.5281/zenodo.11396520 and https://github.com/lseventeen/DIAS.
 Index Terms:
@@
 Owing to its inherently superior spatial and temporal resolution, DSA can accurately reveal lesion details in instances where Computed Tomography Angiography (CTA) and Magnetic Resonance Angiography (MRA) may not offer a definitive diagnosis [2].
-Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
+Thus, DSA is universally acknowledged as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
 Conventionally, neuroradiologists analyze DSA images visually for diagnosis, clinical decision-making, and postoperative evaluation, which is inherently time-consuming, labor-intensive, and subjective.
@@
 Despite its critical importance, cerebrovascular segmentation in DSA has not been extensively researched.
-A key factor contributing to this is the lack of a dedicated DSA dataset specifically designed for developing and evaluating segmentation techniques.
+A contributing factor to this issue is the lack of a dedicated dataset designed for the development and evaluation of vessel segmentation algorithms in DSA.
 However, the creation of such a dataset is particularly challenging due to the unique characteristics of DSA imaging.
@@
 Annotating vessels in DSA necessitates neuroradiologists to meticulously examine each image frame, a process that is both labor-intensive and costly [8].
-Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast, and complex background noise, significantly increases the difficulty of accurate annotation.
-Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8].
-However, these studies rely on in-house datasets, lacking a unified standard for evaluation.
-Most of these menthods [9, 10] focus on segmenting single-frame images, thereby ignoring DSA sequences’ inherent temporal attributes.
-As a result, single-frame based segmentation approaches may not fully represent the intricate vascular morphology of IA, potentially compromising their effectiveness in the diagnosis and treatment of cerebrovascular diseases [8].
-Vessel segmentation in DSA is considered a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map.
+Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast in fine vessels, and complex background noise, significantly increases the difficulty of accurate annotation.
+Currently, several studies [9, 10, 8, 11] have emerged in the domain of vessel segmentation within DSA, demonstrating exceptional performance on in-house datasets.
+Nevertheless, the majority of these methods [9, 10, 11] focus on segmenting single-frame images, thereby neglecting the inherent temporal attributes of DSA sequences.
+Constrained by the fact that single-frame DSA captures only a fraction of the vascular contrast, such methods may inadequately represent the complex vascular structure of IA.
+This shortfall could potentially undermine their utility in the accurate diagnosis and treatment of cerebrovascular diseases [8].
+Vessel segmentation in DSA should be perceived as a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map.
 This approach deviates from the traditional 2D and 3D vessel segmentation methods [12, 13, 14].
@@
 The red arrow in the upper left image indicates segments C1-C3 of the intracranial internal carotid artery (ICA), while the lower left image represents the anterior cerebral artery (ACA) and the middle cerebral artery (MCA).
-In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons.
-Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation.
-To comprehensively evaluate the efficacy of prevalent segmentation methods like fully-supervised, weakly-supervised, and semi-supervised segmentation on the DIAS dataset, we have established a benchmark comprising a vessel sequence segmentation network, a scribble-based learning segmentation framework, and a random patch-based self-training semi-supervised segmentation pipeline.
-Extensive experiments with the DIAS set robust baselines, potentially advancing future algorithmic developments in cerebral vessel segmentation.
-TABLE I: Summary of several public vessel segmentation datasets. ${}^{\dagger}$ represents only training set data is available.
+TABLE I: Summary of several public vessel segmentation datasets. represents only training set data is available.
 Dataset Year Target Modality Dimension Numbers Resolution
@@
 DCA1 [21] 2019 coronary artery angiograpy 2D 134 300$\times$300
-${}^{\dagger}$ [21] 2022 coronary artery CTA 3D 40 512$\times$512$\times$200
-${}^{\dagger}$ [22] 2014 lung vessel CT 3D 20 -
+ASOCA [22] 2022 coronary artery CTA 3D 40 512$\times$512$\times$200
+VESSEL12 [23] 2014 lung vessel CT 3D 20 -
 DIAS 2023 intracranial artery DSA 2D+Time 120 (753) 800$\times$800$\times$4-14
+In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons.
+Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation.
+To comprehensively evaluate the efficacy of prevalent segmentation methods, including fully-supervised, weakly-supervised, and semi-supervised techniques, on the DIAS dataset, we have established a benchmark comprising three key components:
+  • We propose the vessel sequence segmentation network (VSS-Net), incorporating the sequence feature extraction module adept at efficiently capturing spatiotemporal representations of intravascular contrast.
+Notably, VSS-Net is designed to be compatible with any encoder-decoder network, enabling precise segmentation of intracranial arteries in 2D+Time DSA sequences.
+  • We propose an innovative scribble learning-based image segmentation framework, specifically designed for weakly-supervised IA segmentation.
+This framework, distinguished by its dual-model architecture, employs cross pseudo-supervision and consistency regularization, guided by scribble labels.
+This integration significantly boosts the performance of the segmentation network, capitalizing on the synergistic advantages offered by the dual-model approach.
+  • We introduce the random patch-based self-training algorithm, which alleviates the challenges arising from the limited availability of annotated DSA data in IA segmentation by enhancing the model’s generalization through joint training on randomly selected image patches from both labeled and pseudo-labeled data.
+Through extensive experimentation on the DIAS dataset, we have established robust baselines that hold the potential to advance future algorithmic developments in cerebral vessel segmentation.
 II Related work
@@
 Beyond network architectures, some research has focused on optimizing training strategies [36, 37] and modifying loss functions [38].
-However, there have been few studies on DSA vessel segmentation.
-Meng et al.
+There have been few studies on DSA vessel segmentation.
 [9] proposed a CNN-based multiscale dense CNN segmentation network for IA segmentation in single-frame DSA, utilizing multiscale atrous convolution, enhanced dense blocks, and redesigned skip connections.
-Zhang et al.
 [24] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation.
 These two approaches are tailored for vessel segmentation in individual DSA frames.
-For vessel sequence, Hao et al.
-proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features [38].
-Su et al.
-proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8].
+For vessel sequence, [39] proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features.
+[8] proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences.
 While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
 Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers.
-Weakly-supervised segmentation (WSS) trains models with sparse annotations instead of full ones.
+Weakly-Supervised Segmentation (WSS) trains models with sparse or noisy annotations, instead of exact and complete annotations.
 The common methods involves online labeling [40] via consistency checks with the mean teacher or using pseudo-labels supervision [41] derived from model predictions.
+Furthermore, investigating the method for weakly-supervised annotation of vessels in DSA is a research direction deserving of exploration.
+[11] propose a novel learning framework, which utilizes an active contour model for weak supervision and low-cost human-in-the-loop strategies to improve weak label quality.
 Semi-Supervised Segmentation (SSS) aims to explore tremendous unlabeled data with supervision from limited labeled data.
@@
 Furthermore, methods employing consistency regularization strategies [42, 44, 45] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image.
-To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences.
+Various domain adaptation techniques have been proposed to enhance the generalizability of vessel segmentation, leveraging vessel data from other modalities with similar anatomical structures to aid in training, thereby improving the performance of models in SSS.
+[46] propose contrastive SSS for cross anatomy domain adaptation that adapts a model to segment coronary arteries, which requires only limited annotations in the coronary arteries by leveraging existing annotated retinal vessels of similar structure.
+[47] propose a semi-supervised domain adaptation framework for accurately segmenting 3D brain vessels by employing the StyleGAN2 [48] to facilitate the representation of heterogeneous volumetric data and bridge the substantial domain gap between angiography and venography brain images.
+To our knowledge, there is no existing literature on weakly-supervised and semi-supervised vessel segmentation in DSA sequences.
 Building upon our created dataset, we have conducted in-depth research and proposed innovative methods in three distinct areas: network models, WSS, and SSS.
 These approaches are specifically tailored to leverage the unique features of the dataset, addressing each segmentation paradigm comprehensively.
+Furthermore, large-scale pretrained models, i.e., foundation models, exhibit considerable generalizability and adaptability, positioning them as a focal point of recent medical image analysis research [49].
+However, generally, a unified foundation model approach cannot achieve state-of-the-art performance in many medical image analysis tasks due to large variations present in organs and important structures, texture, shape, size and topology (e.g., blood vessels), and imaging modalities [50].
+We are convinced that, through continuous research, foundation models possess the potential to achieve vascular segmentation in DSA that can compete with fully supervised methods, leveraging few-shot learning, zero-shot learning, or prompt engineering.
 III DIAS Dataset
@@
 It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view or lateral view.
-In this study, we retained 120 sequences (753 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates.
+In this study, we retained 120 sequences (762 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates.
 A total of 120 sequences were derived from 60 patients, among whom 24 were female and 36 were male.
 The average age of these patients was 62.5 years, with an age range from 40 to 88 years old.
-The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS).
+The number of frames varies between sequences (6 - 46 frames), but the frame rate is fixed (4 FPS).
 The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt.
@@
 As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
+Figure 2: Full annotated process of DIAS.
 III-B Reference Annotation Protocol
 III-B1 Full Annotation
-Figure 2: Full annotated process of DIAS.
 The manual annotation of IA in DSA sequences necessitates a thorough consideration of all image data within the sequence.
@@
 1 illustrates a typical representation of intracranial vessels across five successive frames during the arterial phase in a lateral view.
-During this phase, the contrast agent progressively spreads from the internal carotid artery (ICA) into cerebral vessels like the anterior cerebral artery (ACA) middle cerebral artery (MCA), carried by the blood flow.
+During this phase, the contrast agent progressively spreads from the Internal Carotid Artery (ICA) into cerebral vessels like the Anterior Cerebral Artery (ACA) and Middle Cerebral Artery (MCA), carried by the blood flow.
 For a single image, the contrast agent is discernible in only a portion of the IA.
-For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+3, it filled the ACA and MCA.
+For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+5, it filled the ACA and MCA.
 To annotate the IA effectively, each image in the sequence must be individually annotated, and then these annotations should be integrated to leverage the full vessel information available in the sequence, thereby ensuring precision in annotation.
@@
 Red represents annotated vessel pixels, blue represents annotated background pixels.
-  • [leftmargin=*,itemsep=-4pt]
   • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the posterior phase.
@@
 RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations.
-In SALE and RDFA, we set the pixel values of the background and vessel annotations to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.
+In SALE and RDFA, we set the pixel values of the background annotation and vessel annotation to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.
 IV Methods
 Automatic IA segmentation in DSA sequences is an under-explored and challenging task.
-To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
+To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using Fully-Supervised Segmentation (FSS), WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
 IV-A Fully-supervised Segmentation
@@
 For IA segmentation in DSA sequence, $M=3$, $N=2$.
-Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA.
+Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA.
 Figure 5 (a,b) illustrates an overview of the proposed network architecture.
-VSS-Net comprises multi-scale sequence feature extraction modules (SFEM).
+In VSS-Net, we employ Sequence Feature Extraction Modules (SFEM) at each stage in the encoder to extract feature maps across multiple scales.
 These modules handle the 2D+Time input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations.
+With the deepening of stages, the input resolution is reduced, consequently expanding the receptive field that facilitates the generation of high-dimensional spatiotemporal representations.
 Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
@@
 Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation.
-Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels.
-This process is formulated as:
-$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i}),$ (1)
-where $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence.
-$W$ are $1\times 1$ convolutional kernels.
-$\ast$ denotes the convolution operation.
-$\oplus$ denotes the concatenation operation.
+Each image in the sequence is concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels.
+The calculation process for the fusion output $Q^{j}$ at scale $j$ is as follows:
+$Q^{j}=\sum_{i=1}^{N}W\ast(x^{j}_{i}\oplus f^{j}_{i}),\quad\text{for }j\in\left\{\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{2^{n-1}}\right\},$ (1)
+where $x^{j}_{i}$ represents the $i^{th}$ frame of the input sequence, scaled to $j$ times the original resolution, $f$ is the feature map, $N$ denotes the length of the sequence, $W$ are $1\times 1$ convolutional kernels, $\ast$ denotes the convolution operation, $\oplus$ denotes the concatenation operation, $n$ denotes the number of stages in the encoder.
 Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of the sequence.
@@
 $h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t},$ (5)
-where $\sigma$ represents the sigmoid function.
-$\tanh$ represents the hyperbolic tangent function.
-$\odot$ denotes the Hadamard product, i.e., element-wise multiplication.
-$W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
+where $\sigma$ represents the sigmoid function, $\tanh$ represents the hyperbolic tangent function, $\odot$ denotes the Hadamard product, i.e., element-wise multiplication, $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
 We simultaneously capture the spatiotemporal dynamics of vascular flow in both forward and backward directions.
@@
 To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels.
-To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization (CR).
+To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization.
 Fig.
 5 (c) presents the implementation framework for SSCR.
-SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
+SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing the segmentation network using mutual consistency and cross pseudo supervision.
 We define $P$ as the set of image pixels and $S\subseteq P$ as the set of annotated pixels.
@@
 The training objective contains three losses: scribble supervision loss $\mathcal{L}_{scr}$, cross pseudo supervision loss $\mathcal{L}_{cps}$, and consistency loss $\mathcal{L}_{con}$:
-$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL)),$ (11)
-$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A})),$ (12)
+$\mathcal{L}_{scr}=$ $\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL),$ (11)
+$\mathcal{L}_{cps}=$ $\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A}),$ (12)
 $\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B}),$ (13)
-where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function.
-$P_{A}$ and $P_{B}$ are the outputs of two models.
-$PL_{A}$ and $PL_{B}$ are the pseudo labels.
-$SL$ is the scribble label.
+where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function, $P_{A}$ and $P_{B}$ are the outputs of two models, $PL_{A}$ and $PL_{B}$ are the pseudo labels, $SL$ is the scribble label.
 Finally, the training objective $\mathcal{L}$ is formulated as:
@@
 Self-training can be simplified into the following steps:
-  1. [leftmargin=*,itemsep=-4pt]
   1. 1.
@@
 V-A Implementation Details and Evaluation Metrics
-TABLE II: Results of fully-supervised segmentation with state-of-the-art models.
-Methods Params / M Metrics
-[…]
-2D UNet [27] 31.04 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
-Res-UNet [51] 9.37 0.7656 0.9655 0.7425 0.9850 0.6235 0.9816 60.56
-UNet++ [52] 9.05 0.7768 0.9659 0.7726 0.9830 0.6374 0.9830 54.65
-Att-UNet [53] 8.73 0.7760 0.9660 0.7686 0.9835 0.6258 0.9811 59.37
-[…]
-MAA-Net [13] 5.80 0.7693 0.9661 0.7423 0.9855 0.6361 0.9818 68.18
-FR-UNet [12] 7.37 0.7779 0.9665 0.7646 0.9841 0.6376 0.9827 53.92
-3D UNet [27] 90.27 0.7691 0.9648 0.7690 0.9824 0.6275 0.9819 62.92
-Res-UNet [51] 26.38 0.7701 0.9653 0.7597 0.9835 0.6284 0.9817 73.05
-UNet++ [52] 26.17 0.7649 0.9639 0.7711 0.9811 0.6214 0.9811 67.02
-Att-UNet [53] 23.6 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
-[…]
-FR-UNet [12] 20.57 0.7635 0.9646 0.7518 0.9830 0.6210 0.9815 65.67
-[…]
-PSC [47] 6.43 0.7649 0.9646 0.7552 0.9832 0.6223 0.9814 76.04
-SVS-Net [38] 8.48 0.7698 0.9659 0.7474 0.9853 0.6277 0.9803 97.28
-ST-UNet [8] 104.37 0.7818 0.9675 0.7533 0.9864 0.6436 0.9843 62.59
-VSS-Net 18.69 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
 The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level.
@@
 We calculated the Dice Similarity Coefficient (DSC), the Area Under the receiver operating Characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and Intersection Over Union (IOU) to evaluate the performance of DSA sequence segmentation.
-Moreover, we introduced the vascular connectivity (VC) [12] metric, which calculates the ratio of connected components in the segmentation image to those in the ground truth, as a means of measuring the connectivity of vessels.
+Furthermore, we conducted evaluations of the Vascular Connectivity (VC) metric [13] and the connectivity-aware similarity measure (clDice) [58].
 V-B Fully-supervised Segmentation
@@
 $y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x))).$ (16)
-Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods.
-Using FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe.
-Specifically, Sen has increased by 1.3%.
+[…]
+Methods Params / M Metrics
+[…]
+2D UNet 31.04 0.7700$\pm$0.0525 0.9654$\pm$0.0125 0.7570$\pm$0.0879 0.9839$\pm$0.0098 0.6290$\pm$0.0695 0.9829$\pm$0.0080 0.6939$\pm$0.0628 63.60$\pm$39.01
+Res-UNet 9.37 0.7656$\pm$0.0548 0.9655$\pm$0.0120 0.7425$\pm$0.0950 0.9850$\pm$0.0083 0.6235$\pm$0.0719 0.9816$\pm$0.0093 0.6870$\pm$0.0745 60.51$\pm$40.00
+UNet++ 9.05 0.7767$\pm$0.0468 0.9659$\pm$0.0122 0.7726$\pm$0.0752 0.9830$\pm$0.0091 0.6374$\pm$0.0630 0.9830$\pm$0.0084 0.7042$\pm$0.0589 54.79$\pm$32.24
+Att-UNet 8.73 0.7760$\pm$0.0435 0.9660$\pm$0.0118 0.7686$\pm$0.0800 0.9835$\pm$0.0092 0.6361$\pm$0.0587 0.9837$\pm$0.0077 0.7030$\pm$0.0569 59.39$\pm$31.50
+CS2-Net 8.93 0.7747$\pm$0.0436 0.9664$\pm$0.0120 0.7524$\pm$0.0878 0.9855$\pm$0.0078 0.6344$\pm$0.0586 0.9833$\pm$0.0083 0.6941$\pm$0.0588 64.91$\pm$45.90
+MAA-Net 5.80 0.7693$\pm$0.0508 0.9661$\pm$0.0120 0.7423$\pm$0.0969 0.9855$\pm$0.0091 0.6279$\pm$0.0669 0.9818$\pm$0.0095 0.69$\pm$0.0690 68.20$\pm$42.60
+FR-UNet 7.37 0.7779$\pm$0.0459 0.9665$\pm$0.0118 0.7646$\pm$0.0808 0.9841$\pm$0.0083 0.6376$\pm$0.0615 0.9827$\pm$0.0078 0.7112$\pm$0.0603 53.92$\pm$27.93
+3D UNet 90.27 0.7692$\pm$0.0490 0.9648$\pm$0.0121 0.7690$\pm$0.0887 0.9824$\pm$0.0090 0.6275$\pm$0.0650 0.9818$\pm$0.0090 0.6975$\pm$0.0663 62.95$\pm$37.21
+Res-UNet 26.38 0.7701$\pm$0.0454 0.9653$\pm$0.0121 0.7597$\pm$0.0809 0.9835$\pm$0.0087 0.6284$\pm$0.0606 0.9817$\pm$0.0092 0.6991$\pm$0.0606 73.19$\pm$45.50
+UNet++ 26.17 0.7649$\pm$0.0448 0.9639$\pm$0.0121 0.7712$\pm$0.0791 0.9811$\pm$0.0088 0.6215$\pm$0.0589 0.9811$\pm$0.0092 0.6989$\pm$0.0595 67.03$\pm$48.51
+Att-UNet 23.60 0.7677$\pm$0.0513 0.9647$\pm$0.0131 0.7570$\pm$0.0816 0.9829$\pm$0.0106 0.6258$\pm$0.0686 0.9811$\pm$0.0098 0.6966$\pm$0.0624 68.32$\pm$40.96
+CS2-Net 23.83 0.7687$\pm$0.0472 0.9645$\pm$0.0124 0.7712$\pm$0.0838 0.9816$\pm$0.0099 0.6267$\pm$0.0628 0.9808$\pm$0.0100 0.6960$\pm$0.0650 67.91$\pm$45.25
+FR-UNet 20.57 0.7635$\pm$0.0587 0.9646$\pm$0.0123 0.7518$\pm$0.0889 0.9830$\pm$0.0107 0.6211$\pm$0.0754 0.9815$\pm$0.0096 0.6913$\pm$0.0686 65.65$\pm$45.09
+[…]
+PSC 6.43 0.7649$\pm$0.0523 0.9646$\pm$0.0123 0.7552$\pm$0.0896 0.9832$\pm$0.0095 0.6223$\pm$0.069 0.9814$\pm$0.0104 0.6965$\pm$0.0604 76.05$\pm$44.63
+SVS-Net 8.48 0.7697$\pm$0.0435 0.9659$\pm$0.0115 0.7474$\pm$0.0826 0.9853$\pm$0.0072 0.6277$\pm$0.0578 0.9803$\pm$0.0085 0.6880$\pm$0.0580 97.33$\pm$49.10
+ST-UNet 104.37 0.7818$\pm$0.0406 0.9675$\pm$0.0119 0.7533$\pm$0.0821 0.9864$\pm$0.0070 0.6435$\pm$0.0551 0.9843$\pm$0.0075 0.7073$\pm$0.0587 62.54$\pm$38.25
+VSS-Net 18.69 0.7822$\pm$0.0419 0.9668 $\pm$ 0.0119 0.7770$\pm$0.0830 0.9838$\pm$0.0076 0.6442$\pm$0.0568 0.9857$\pm$0.0070 0.7119$\pm$0.0583 43.02$\pm$27.47
+Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC, clDice and VC, among all the evaluated methods.
+Using 2D FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe.
+Specifically, Sen has increased by 1.24%.
 ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe.
 However, the number of model parameters is approximately 5.58 times that of VSS-Net.
-The VC score that is 19.57 higher than VSS-Net indicates poorer vascular connectivity.
+The VC score that is 19.52 higher than VSS-Net indicates poorer vascular connectivity.
 Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages.
 PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures.
+Figure 6: Dimensionality reduction operation applied to 2D/3D models
 Figure 7: Visualization of segmentation results of state-of-the-art models.
@@
 We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Fig.
-6.
+7.
 In the figure, the red and blue colors represent pixels that are false negatives and false positives, respectively, indicative of segmentation errors.
@@
 Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Fig.
-6.
+7.
 In comparison with other methods, VSS-Net has the highest number of white pixels in the fine vessels, indicating the best segmentation results.
@@
 […]
-pCE [47] 0.7455 0.9586 0.7845 0.9735 0.5966 0.9379 60.11 0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
-EM [56] 0.7458 0.9575 0.8029 0.9707 0.5970 0.9725 41.18 0.6356 0.9226 0.8508 0.9281 0.4699 0.9618 9.05
-RLoss [45] 0.7146 0.9600 0.6539 0.9867 0.5581 0.9367 46.30 0.5916 0.9147 0.7812 0.9257 0.4232 0.9475 5.17
-IVM [46] 0.6606 0.9299 0.8546 0.9360 0.4971 0.9455 71.58 0.6322 0.9198 0.8760 0.9232 0.4662 0.9628 14.78
-DMPLS [40] 0.7471 0.9581 0.7963 0.9717 0.5995 0.9708 37.56 0.6327 0.9208 0.8508 0.9259 0.4668 0.9643 9.86
-SSCR 0.7589 0.9603 0.8079 0.9732 0.6141 0.9761 36.52 0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
+pCE 0.7455 0.9586 0.7845 0.9735 0.5966 0.9379 0.7022 60.11 0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 0.6589 10.02
+EM 0.7458 0.9575 0.8029 0.9707 0.5970 0.9725 0.7103 41.18 0.6356 0.9226 0.8508 0.9281 0.4699 0.9618 0.6576 9.05
+RLoss 0.7146 0.9600 0.6539 0.9867 0.5581 0.9367 0.5906 46.30 0.5916 0.9147 0.7812 0.9257 0.4232 0.9475 0.5702 5.17
+IVM 0.6606 0.9299 0.8546 0.9360 0.4971 0.9455 0.6910 71.58 0.6322 0.9198 0.8760 0.9232 0.4662 0.9628 0.6738 14.78
+DMPLS 0.7471 0.9581 0.7963 0.9717 0.5995 0.9708 0.6996 37.56 0.6327 0.9208 0.8508 0.9259 0.4668 0.9643 0.6601 9.86
+SSCR 0.7651 0.9639 0.7620 0.9810 0.6214 0.9778 0.6977 34.48 0.6574 0.9277 0.8398 0.9344 0.4958 0.9672 0.6732 6.47
 Figure 8: Visualization of the segmentation results of scribble supervised segmentation methods trained on RDFA and SALE annotations.
@@
 Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods.
-It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 4.67% higher compared to SSCR, and it is 4.12% higher on SALE.
+It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 9.26% higher compared to SSCR, and it is 3.62% higher on SALE.
 Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background.
@@
 […]
-0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 10.02
-0.6364 0.9244 0.8346 0.9316 0.4706 0.9479 6.57
-0.6472 0.9275 0.8385 0.9346 0.4827 0.9445 7.29
-0.6484 0.9276 0.8536 0.9334 0.4827 0.9639 8.28
-0.6722 0.9343 0.8348 0.9421 0.5115 0.9658 10.30
-TABLE V: Ablation research results of RPST for semi-supervised IA sequence segmentation.
+0.6396 0.9246 0.8462 0.9308 0.4738 0.9560 0.6589 10.02
+0.6364 0.9244 0.8346 0.9316 0.4706 0.9479 0.6437 6.57
+0.6472 0.9275 0.8385 0.9346 0.4827 0.9445 0.6692 7.29
+0.6484 0.9276 0.8536 0.9334 0.4827 0.9639 0.6685 8.28
+0.6574 0.9277 0.8398 0.9344 0.4958 0.9672 0.6732 6.47
+TABLE V: Ablation research results of RPST for semi-supervised IA sequence segmentation. (L: Labeled data, U: Unlabeled data)
 Methods Data Metrics
 […]
-Teacher 1 0 0.6266 0.9259 0.6853 0.9451 0.4693 0.8538 385.62
-Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 107.35
-S w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 108.96
-Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 103.09
-Teacher 3 0 0.6919 0.9490 0.7358 0.9673 0.5336 0.9182 320.02
-Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 121.84
-S w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.5700 0.9578 125.83
-Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 125.78
-Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 148.89
-Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 93.31
-S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
-Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 87.72
+Teacher 1 0 0.6266 0.9259 0.6853 0.9451 0.4693 0.8538 0.5173 385.62
+Student 1 30 0.6486 0.9294 0.7085 0.9464 0.4941 0.9505 0.5668 107.35
+S w/o SDA 1 60 0.6394 0.9296 0.6922 0.9483 0.4832 0.9352 0.5531 108.96
+Student 1 60 0.6502 0.9305 0.7177 0.9482 0.4958 0.9433 0.5653 103.09
+Teacher 3 0 0.6919 0.9490 0.7358 0.9673 0.5336 0.9182 0.5921 320.02
+Student 3 30 0.7265 0.9555 0.7478 0.9721 0.5745 0.9591 0.6523 121.84
+S w/o SDA 3 60 0.7228 0.9555 0.7456 0.9734 0.5700 0.9578 0.6404 125.83
+Student 3 60 0.7275 0.9549 0.7595 0.9778 0.5893 0.9616 0.6502 125.78
+Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 0.6512 148.89
+Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 0.6747 93.31
+S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 0.6708 93.92
+Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 0.6794 87.72
 We conducted ablation studies to investigate the impact of various components in our proposed framework for weakly-supervised IA segmentation.
 As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis.
-Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as DSC and AUC.
+Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as Sen, AUC and clDice.
 Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC.
@@
 We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models.
-The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 2.5% and 2.38%, respectively.
+The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 1.02% and 0.9%, respectively.
 V-D Semi-supervised Segmentation
@@
 When using only the first image as the model input, the large vessels are accurately segmented.
-However, due to the absence of angiographic information for fine vessels in the first image frame, the thin vessel at the arterial termini within the middle cerebral artery (MCA) remain incomplete.
+However, due to the absence of angiographic information for fine vessels in the first image frame, the thin vessel at the arterial termini within the MCA remain incomplete.
 Conversely, using only the last image as input enables the accurate segmentation of most small vessels, yet it falls short in precisely delineating the larger vessels in proximity to the ICA.
@@
 […]
-First image 0.6232 0.9515 0.5388 0.9864 0.4601 0.9049 12.05
-Last image 0.7604 0.9647 0.7394 0.9842 0.6146 0.9813 51.89
-First + Last image 0.7777 0.9668 0.7589 0.9850 0.6373 0.9846 44.22
-Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 39.25
-Min 0.7802 0.9673 0.7631 0.9854 0.6418 0.9842 40.02
-Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
+First 0.6232 0.9515 0.5388 0.9864 0.4601 0.9049 0.6835 12.05
+Last 0.7604 0.9647 0.7394 0.9842 0.6146 0.9813 0.6940 51.89
+First+Last 0.7777 0.9668 0.7589 0.9850 0.6373 0.9846 0.7039 44.22
+Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 0.6997 39.25
+Min 0.7802 0.9673 0.7631 0.9854 0.6418 0.9842 0.7040 40.02
+Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 0.7119 43.02
 The combination of early and late arterial angiography, known as ‘First+Last’, yielded favorable results, surpassing those of both inputs used independently.
 As shown in Fig.
-9 (a) and (b), although the combination of the ’First’ and ’Last’ images from the arterial phase can display the majority of the vascular structures filled with contrast agent, we can still utilize the intermediate images to supplement details.
+10 (a) and (b), although the combination of the ‘First’ and ‘Last’ images from the arterial phase can display the majority of the vascular structures filled with contrast agent, we can still utilize the intermediate images to supplement details.
 The experimental findings indicate that employing the complete DSA sequence as input is imperative for achieving optimal performance.
@@
 In this paper, we present the DSA-sequence IA segmentation dataset, DIAS, which provides pixel-level annotations along with two forms of weakly supervised annotations.
-We establish a benchmark for evaluating IA sequence segmentation on DIAS, comprising (1) the Vessel Sequence Segmentation Network, VSS-Net; (2) the Scribble Learning-based Segmentation Framework, SSCR; and (3) the Random Patch-based Self-Training Framework, RPST.
+We establish a benchmark for evaluating IA sequence segmentation on DIAS dataset, comprising (1) the vessel sequence segmentation network, adept at capturing the spatiotemporal representations in DSA angiography; (2) the weakly-supervised segmentation framework integrating scribble-supervision with consistency regularization; and (3) the random patch-based self-training framework, strategically designed to utilize unannotated DSA data to enhance segmentation performance.
 Leveraging the DIAS dataset, we have conducted comprehensive experiments to validate the effectiveness of our proposed methodologies, facilitating a deeper understanding of their performance characteristics and offering valuable insights for future research endeavors.
-Looking ahead, our research will persist, including the expansion of segmentation targets to venous and the segmentation of arteries from complete angiographic sequences.
-We emphasize that the DIAS dataset and benchmark have been made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.
+Looking ahead, our research will continue, including the expansion of segmentation targets to veins and the segmentation of arteries from complete angiographic sequences.
+We would like to emphasize that the DIAS dataset and its associated benchmark will be made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.